### v1

# Pushing the Complexity Boundaries of Fixed-Point Equations: Adaptation to Contraction and Controlled Expansion

Abstract
Fixed-point equations with Lipschitz operators have been studied for more than a century, and are central to problems in mathematical optimization, game theory, economics, and dynamical systems, among others. When the Lipschitz constant of the operator is larger than one (i.e., when the operator is expansive), it is well known that approximating fixed-point equations becomes computationally intractable even in basic finite-dimensional settings. In this work, we aim to push these complexity boundaries by introducing algorithms that can address problems with mildly expansive (i.e., with Lipschitz constant slightly larger than one) operators not excluded by existing lower bounds, attaining the best possible fixed-point error up to universal constants. We further introduce a class of gradually expansive operators that allow for constant (up to $\approx 1.4$) expansion between points, for which we prove convergence to $\epsilon$-approximate fixed points in order-$(1/\epsilon)$ iterations for $\epsilon>0.$ Our algorithms automatically adapt to the Lipschitz constant of the operator and attain optimal oracle complexity bounds when the input operator is nonexpansive or contractive. Our results apply to general, possibly infinite-dimensional normed vector spaces and can be extended to positively curved geodesic metric spaces.
1 Introduction
Fixed-point operator equations in a normed vector space $({\mathcal{E}},\left\lVert\cdot\right\rVert)$ ask for identifying an element of the space ${\bm{x}}\in{\mathcal{E}}$ with the property that ${\bm{T}}({\bm{x}})={\bm{x}},$ where ${\bm{T}}:{\mathcal{E}}\to{\mathcal{E}}$ is a given operator (or a mapping). If there exists $\gamma\geq 0$ such that for any ${\bm{x}},{\bm{y}}\in{\mathcal{E}},$ we have $\left\lVert{\bm{T}}({\bm{x}})-{\bm{T}}({\bm{y}})\right\rVert\leq\gamma\left\lVert{\bm{x}}-{\bm{y}}\right\rVert,$ then ${\bm{T}}$ is said to be $\gamma$-Lipschitz. While simple to state, fixed-point equations encompass a broad range of problems in engineering, game theory, economics, mathematical optimization, and dynamical systems, among others (see, for instance, [FP03, Yan09, Tod13] for many illustrative examples across these areas). Given their centrality in these scientific disciplines, the topic of how to efficiently solve fixed-point equations has been intensely investigated over the past several decades.
Classical results in this domain have provided iterative algorithms based on oracle access to ${\bm{T}}$ (i.e., algorithms that only rely on evaluations of ${\bm{T}}$) with guaranteed convergence whenever the operator is contractive (i.e., Lipschitz-continuous with $\gamma<1$), nonexpansive (i.e., Lipschitz-continuous with $\gamma=1$), or satisfies other slightly more general but closely related properties such as quasi-nonexpansiveness or pseudo-contractivity; see [Bro67, Ish74, Man53, Kra55, Hal67, Ede66, Opi67]. Such results have been established in broad generality, even applying to possibly infinite-dimensional normed vector spaces. At present, the complexity of solving fixed-point equations for contractive and nonexpansive operators (i.e., with Lipschitz constant $\gamma\leq 1$) is well-understood [Cho87, Nem91, BB96, Sik09, SS17, Lie21, Kim21, DW22, PR22, BCPS19, BC24], with some of the recent results even characterizing the complexity of common algorithms down to constants [BCC21, CC23]. For operators that are allowed even a small amount of expansion (i.e., with Lipschitz constant slightly larger than one), existing results require a type of a (local) error bound that relates the fixed-point error $\|{\bm{T}}({\bm{x}})-{\bm{x}}\|$ to some notion of distance between ${\bm{x}}$ and a nearest fixed point to establish nonasymptotic (local or global) convergence of considered algorithms [Nat93, LTT18, LL21, HLS23].
[…]
The above discussion raises the question:
What kind of problems lie between the efficiently solvable fixed-point problems with nonexpansive operators and the notoriously hard problems with general Lipschitz operators?
Put differently, what are some general conditions for Lipschitz-continuous operators ${\bm{T}}:{\mathcal{E}}\to{\mathcal{E}}$ under which efficient convergence to approximate fixed points can be guaranteed? In addition to being a well-recognized research challenge [BB96, Das22], this question is not merely of theoretical interest: such problems commonly arise in a variety of applications like machine learning (e.g., in training of Generative Adversarial Networks or in Multi-Agent Reinforcement Learning; see [Das22]), feasibility problems and structured nonconvex optimization [LTT18], or in X-ray free electron laser (X-FEL) imaging [HLS23]. A desirable property of algorithms addressing such problems is that they can gracefully adapt to input instances, in the sense that if an input instance happens to be defined by a nonexpansive or a contractive operator, the algorithm exhibits similar convergence properties to optimal algorithms for the corresponding class of instances. The primary contributions of this work lie in this general domain.
1.1 Contributions and Paper Overview
Our work starts from investigating the properties of a simple variant of classical Halpern iteration [Hal67], which fixes the step size $\lambda_{k}$ in the classical iteration ${\bm{x}}_{k+1}=\lambda_{k}{\bm{x}}_{0}+(1-\lambda_{k}){\bm{T}}({\bm{x}}_{k})$ to some $\lambda\in(0,1);$ namely, the iteration is:
${\bm{x}}_{k+1}=\lambda{\bm{x}}_{0}+(1-\lambda){\bm{T}}({\bm{x}}_{k}).$ (FixHal)
The usefulness of considering such an iteration is not clear a priori, as it cannot be guaranteed to converge to zero fixed point error $\|{\bm{T}}({\bm{x}}_{k})-{\bm{x}}_{k}\|,$ even asymptotically (as $k\to\infty$) and assuming the operator ${\bm{T}}$ is contractive. [^2] Consider, for instance, ${\bm{T}}({\bm{x}})=\gamma{\bm{x}}$ for $\gamma\in(0,1)$. Then ${\bm{x}}_{k}\to\frac{\lambda}{1-(1-\lambda)\gamma}{\bm{x}}_{0}$ as $k$ tends to infinity, while the only fixed point of this operator is the zero element of the space. Thus, unless initialized at the fixed point, (FixHal) does not converge to it. However, as we show in Lemma 1, for any target error $\epsilon>0$ and a $\gamma$-Lipschitz operator with $\gamma\in(0,1],$ it is possible to choose a value of the step size $\lambda\in(0,1)$ such that (FixHal) converges to a point ${\bm{x}}_{k}$ with fixed-point error $\|{\bm{T}}({\bm{x}}_{k})-{\bm{x}}_{k}\|\leq\epsilon$ within the number of iterations competitive with the number of iterations of algorithms that are optimal (in terms of their query oracle complexity) when the operator is nonexpansive ($\gamma=1$) or contractive ($\gamma<1$).
An appealing property of (FixHal) that we leverage is that it contracts distances between successive iterates of nonexpansive operators, since for any $k\geq 1,$
$\|{\bm{x}}_{k+1}-{\bm{x}}_{k}\|=(1-\lambda)\|{\bm{T}}({\bm{x}}_{k})-{\bm{T}}({\bm{x}}_{k-1})\|.$ (1)
We observe that this property is preserved even if the operator ${\bm{T}}$ is not nonexpansive, but allows for slight expansion, with its Lipschitz constant $\gamma$ being slightly larger than one (e.g., $\gamma=1+\lambda/2$). We use this observation to argue that (FixHal) can be used to approximate fixed points of operators that are mildly expansive (or “almost nonexpansive;” see [LTT18]). In particular, if ${\bm{T}}$ is a $\gamma$-Lipschitz operator that maps a compact convex set of diameter $D$ to itself and if $\gamma<1+\epsilon/D$ for some $\epsilon>0$, then applying (FixHal) with an appropriate choice of the step size $\lambda$ can ensure convergence to a point with a fixed-point error $\epsilon$. Further, the number of required iterations is near-linear in $1/\epsilon$ as long as $\gamma\leq 1+c\epsilon/D$ for any $c\in(0,1)$ such that $1/(1-c)={\mathcal{O}}(1).$ Although this result may not seem impressive on its own, we note that any algorithm that is based solely on oracle queries to ${\bm{T}}$ when applied to a $\gamma$-Lipschitz operator ${\bm{T}}:[0,D]^{d}\to[0,D]^{d}$ with $\gamma\geq 1+C\epsilon/D$ for an absolute constant $C$ (independent of any of the problem parameters), must make an exponential in the dimension number of queries to ${\bm{T}}$ to find ${\bm{x}}$ with fixed-point error $\|{\bm{T}}({\bm{x}})-{\bm{x}}\|\leq\epsilon$ [HPV89]; see Section 2.2 for a more detailed discussion. Thus, although modest, up to an absolute constant $C$, our result gets to the boundary of what is information-theoretically possible.
Starting from these basic results about (FixHal) (which are relatively simple to prove), we then expand our discussion and contributions by introducing algorithms that gradually decrease the value of the step size parameter $\lambda$. We first consider the case in which $D$, (an upper bound on) the diameter of the set containing algorithm iterates, is known, but the operator ${\bm{T}}$ is not necessarily nonexpansive (though it is $\gamma$-Lipschitz). In Section 3, we first introduce Gradual Halpern Algorithm (GHAL, Algorithm 1), which gradually adjusts the step size $\lambda$ as it progresses. Given $\epsilon>0$ and without an a priori knowledge of the Lipschitz constant $\gamma,$ GHAL simultaneously recovers the optimal oracle complexity for both nonexpansive ($\gamma=1$) and contractive ($\gamma<1$) operators, as long as $\gamma$ is not too small [^3] In particular, as long as $\gamma=\Omega(\sqrt{\epsilon/\left\lVert{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\right\rVert})$. (see [TW84, SW87, Cho87, Nem91, DW22, PR22] for lower bounds). Our second algorithm, Adaptive Gradual Halpern Algorithm (AdaGHAL, Algorithm 2), further removes the assumption about knowledge of $D$ for operators that are nonexpansive or contractive (and is thus parameter-free), while maintaining the convergence properties of GHAL.
We also prove approximation error guarantees and convergence bounds for GHAL when $\gamma>1$ and ${\bm{T}}$ maps a convex compact set of diameter $D$ to itself. In particular, we prove that GHAL can guarantee convergence to a point ${\bm{x}}$ with fixed-point error $\left\lVert{\bm{T}}({\bm{x}})-{\bm{x}}\right\rVert\leq c(\gamma-1)D$ within ${\mathcal{O}}(\min\{D/\epsilon,1/(\gamma-1)\})$ iterations, each querying ${\bm{T}}$ a constant number of times, for any $c>1$ such that $1/(c-1)={\mathcal{O}}(1).$ This result is primarily useful for values of $\gamma$ between one and two, since when the diameter of the set is bounded by $D,$ we have that $\|{\bm{T}}({\bm{x}})-{\bm{x}}\|\leq D.$ For $\gamma<1+\epsilon/D,$ this result guarantees convergence to a point with fixed-point error $\epsilon$ and, as argued above for (FixHal), gets to the boundary of what is information-theoretically possible for polynomial-time oracle-based algorithms.
As our final contribution, in Section 4, we introduce a class of $\alpha$-gradually expansive operators, defined as operators mapping a convex compact set ${\mathcal{C}}$ of diameter $D$ to itself with the property that for some $\alpha\in[0,1)$ and all ${\bm{x}},{\bm{y}}\in{\mathcal{C}}$, it holds that
$\|{\bm{T}}({\bm{x}})-{\bm{T}}({\bm{y}})\|\leq\Big{(}1+\frac{\alpha\max\{\left\lVert{\bm{T}}({\bm{x}})-{\bm{x}}\right\rVert,\,\left\lVert{\bm{T}}({\bm{y}})-{\bm{y}}\right\rVert\}}{D}\Big{)}\|{\bm{x}}-{\bm{y}}\|.$ (2)
Observe that such operators allow for the Lipschitz constant of ${\bm{T}}$ to be as large as $1+\alpha$. Moreover, large expansion is possible even between fixed points of ${\bm{T}}$ and other points in the set, due to the “max” of fixed-point errors in the condition (2). We prove that for any $\epsilon>0,$ GHAL applied to an $\alpha$-gradually expansive operator ${\bm{T}}$ with $\alpha\leq 0.4$ outputs a point ${\bm{x}}\in{\mathcal{C}}$ with fixed-point error $\left\lVert{\bm{T}}({\bm{x}})-{\bm{x}}\right\rVert\leq\epsilon$ after at most ${\mathcal{O}}(D/\epsilon)$ iterations, each requiring a constant number of oracle queries to ${\bm{T}}.$
We note that all our results apply to arbitrary normed real vector spaces $({\mathcal{E}},\left\lVert\cdot\right\rVert),$ including those of infinite dimension. Further, they can be generalized to positively curved geodesic metric spaces known as Busemann spaces; see Section 2.3 for a discussion on how this can be done.
Finally, we provide simple numerical examples in Section 5 to illustrate the performance of proposed algorithms. More comprehensive numerical evaluation of the algorithms, although of interest, is beyond the scope of the present work, as our primary contributions are theoretical. We conclude the paper with a discussion of open questions.
1.2 Further Notes on Related Work
As mentioned earlier in the introduction, there is a long line of related work on solving fixed-point equations, spanning areas such as convex analysis, mathematical programming, game theory, and theoretical computer science. Here we focus on reviewing those works that are most closely related to ours.
As mentioned earlier, the basic iteration (FixHal) that underlies our algorithms is a fixed-step variant of classical Halpern iteration [Hal67]. Halpern iteration was originally developed as a method for solving fixed-point equations with nonexpansive operators. Its convergence properties have been studied for many decades. For instance, asymptotic convergence was established by [Wit92], while first nonasymptotic results about convergence of Halpern iteration were established by [Leu07], using proof-theoretic techniques of [Koh08]. These first nonasymptotic convergence results from [Leu07] are rather loose, showing that the fixed point error $\|T({\bm{u}}_{k})-{\bm{u}}_{k}\|$ decays at rate roughly of the order $1/\log(k)$. This convergence rate was tightened to order-$(1/\sqrt{k})$ by [Koh11]. The later results of [SS17] imply that the convergence rate of Halpern iteration with step size $\lambda_{k}=1/(k+1)$ can be tightened to order-$\log(k)/k$, while [CC23] provided tight rate estimates for different step sizes, proving, among other results, that a slightly different step size $\lambda_{k}$ (but still of order $1/k$) leads to the optimal convergence rate of order $1/k.$ All these results [Leu07, Koh11, SS17, CC23] apply to general normed real vector spaces. Further, stochastic variants of Halpern iteration have been used to address problems in reinforcement learning, where the focus is on $\ell_{\infty}$ setups; see, for instance, [BC24, LBC25].
In a somewhat separate line of work, convergence of Halpern iteration was also investigated separately in Euclidean (or, slightly more generally, Hilbert) spaces. Here, the work of [Lie21] and slightly later of [Kim21] provide rates of the order $1/k$ for Halpern iteration with step size $\lambda_{k}=\Theta(1/k)$ (slightly tighter than the result from [SS17]), using a computer-assisted approach of [DT14] to analyze Halpern iteration. In [Dia20], methods based on Halpern iteration were developed to solve monotone inclusion (or root-finding problems) for monotone Lipschitz operators, using equivalence (up to scaling and translation) between solving fixed-point equations with nonexpansive operators and solving root-finding problems for cocoercive operators, which holds in Hilbert spaces (but not more generally). The same work used a restart-based strategy to address root-finding problems with strongly monotone Lipschitz operators, and the basic ideas from this strategy can be used to address fixed-point problems with potentially contractive operators. These results were further extended to remove logarithmic factors by introducing extragradient-like variants of Halpern iteration in [YR21, TDL21, LK21], analyze a variant with adaptive step sizes in [HXDM24], and to address stochastic and finite-sum settings in [CSGD22, CAD24, CL24].
Further, [Dia20] argued that in high dimensions, [^4] For low-dimensional settings with nonexpansive operators, the optimal oracle complexity is of the order $d\log(1/\epsilon),$ where $d$ is the dimension and $\epsilon$ the target fixed-point error. We refer to [Sik09] for a relevant discussion of related work. the convergence rate $1/k$ of Halpern iteration is unimprovable up to a logarithmic factor, using a reduction from the lower bounds for min-max optimization in [OX21]. This lower bound was tightened to remove the logarithmic factor (even for affine operators) in [DW22], where it was also proved that the convergence rate of the classical Krasnoselskii-Mann (KM) iteration [Man53, Kra55] (and similar iterations; see [DW22]) cannot be improved beyond $1/\sqrt{k}.$ Similar results were later obtained independently by [CC23] and [PR22], with [PR22] also improving the constant in the general lower bound $\Omega(1/k)$ from [DW22] for solving fixed-point equations with nonexpansive operators.
Moving beyond nonexpansive operators, as mentioned in the introduction, approximating fixed-point equations quickly becomes intractable. For ${\bm{T}}:[0,1]^{d}\to[0,1]^{d}$ that is $\gamma$-Lipschitz w.r.t. the $\ell_{\infty}$ norm, with $\gamma>1,$ the classical results of [HPV89] lead to exponential in the dimension oracle complexity lower bound for approximating the fixed-point error to some $\epsilon\in(0,1/10)$, already for $\gamma=1+c\epsilon$, where $c$ is an absolute constant (see Section 2.2 for a more detailed discussion). The challenges associated with solving fixed-point equations with general Lipschitz operators have also motivated the introduction of different complexity classes in theoretical computer science; see [Yan09] for an overview. It was recently shown in [ADD+25] that similar to [HPV89] exponential lower bounds also apply to operators mapping the unit $d$-dimensional Euclidean ball to itself and that are $\gamma$-Lipschitz with $\gamma>1$ w.r.t. the $\ell_{2}$ norm. Further, it was shown in the same work that this exponential oracle complexity lower bound cannot be broken by slightly perturbing the problem, in the framework of smoothed analysis [ST04].
We point out here that in these two (finite-dimensional $\ell_{\infty}$ and $\ell_{2}$) settings on convex compact sets, the existence of a fixed point can be guaranteed via classical Brouwer fixed-point theorem [Bro11]. It is well-known that this fixed-point theorem fails in infinite-dimensional Banach spaces even for nonexpansive operators, [^5] Recall here that classical Banach fixed-point theorem [Ban22] that guarantees existence of fixed points requires the operator to be contractive. while for a $\gamma$-Lipschitz operator ${\bm{T}}:{\mathcal{C}}\to{\mathcal{C}}$ with $\gamma>1,$ where ${\mathcal{C}}$ is a convex compact set of diameter $D$, we have $\inf_{{\bm{x}}\in{\mathcal{C}}}\|{\bm{T}}({\bm{x}})-{\bm{x}}\|\leq(1-1/\gamma)D$ and this inequality is tight in general [Goe73]; see also Section 2.2 for a related discussion. An open question regarding finding solutions ${\bm{x}}\in{\mathcal{C}}$ with $\|{\bm{T}}({\bm{x}})-{\bm{x}}\|\leq(1-1/\gamma)D+\epsilon$ for $\epsilon>0$ was raised in [BB96] and remains open. We conjecture that such a result is not possible. Our work leads to a bound of the form $\|{\bm{T}}({\bm{x}})-{\bm{x}}\|\leq(\gamma-1)D+\epsilon$ within ${\mathcal{O}}\big{(}\frac{\gamma\log(\left\lVert{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\right\rVert/\epsilon)}{\epsilon}\big{)}$ queries to ${\bm{T}}$ and is meaningful for $\gamma\in(1,2)$.
Finally, we note that there is a line of work that establishes convergence results for solving fixed-point equations with $\gamma$-Lipschitz operators with $\gamma>1$, assuming there is an appropriate notion of a “local error bound” (see [Pan97] for a classical survey on this topic) that bounds distance to fixed-points of ${\bm{T}}({\bm{x}})$ by its fixed-point error $\left\lVert{\bm{T}}({\bm{x}})-{\bm{x}}\right\rVert$. The first such result we are aware of is [Nat93], which provided an algorithm for solving $\epsilon$-approximate fixed points in the $\ell_{\infty}$ setting discussed above for [HPV89]. The oracle complexity of the algorithm from [Nat93] is logarithmic in $1/\epsilon$, but is however exponential otherwise, scaling with $(2\kappa(\gamma-1))^{d},$ where $\kappa$ is a “condition number” associated with the assumed local error bound, further assumed to be bounded by $1/\epsilon.$ Additionally, a recent line of work [LTT18, LL21, HLS23] provided convergence results for solving fixed-point equations with “almost averaged operators,” under a similar (though relaxed) local error bound known as metric (sub)regularity. Almost averaged operators are defined in [LTT18] as operators ${\bm{T}}$ expressible as ${\bm{T}}=(1-\alpha){\bm{\mathrm{Id}}}+\alpha\widetilde{\bm{T}}$, where $\alpha\in(0,1),$ ${\bm{\mathrm{Id}}}$ is the identity operator, and $\widetilde{\bm{T}}$ is an operator that is $\gamma$-Lipschitz with $\gamma=\sqrt{1+\epsilon}$ for some $\epsilon\in[0,1)$. Under such a set of assumptions and with an appropriate condition that relates the constant of the local error bound to $\epsilon$ and $\alpha,$ it is possible to establish contraction of distance between algorithm iterates and the fixed points of ${\bm{T}},$ leading to (usually local and often linear) convergence. We consider this line of work based on local error bounds complementary to ours.
2 Convergence of the Fixed-Step Halpern Iteration
In this section, we analyze the basic fixed-step Halpern iteration (FixHal). Clearly, because the step size does not diminish as we iterate, we cannot eliminate the contribution of the initial point ${\bm{x}}_{0}$ and thus cannot hope for the iterates to always converge to a vector ${\bm{x}}$ with an arbitrarily small error. Nevertheless, convergence to an $\epsilon$-approximate fixed point is possible for a sufficiently small step size $\lambda(\epsilon)$ and it applies even if the operator is not guaranteed to be nonexpansive (but only mildly expansive), as we argue in this section. Additionally, convergence results from this section will serve as basic ingredients for obtaining more sophisticated results in later sections.
2.1 Convergence for Nonexpansive Operators
The main motivation for considering (FixHal) came from an observation about contraction of distances between successive iterates, as discussed in the introduction. [^6] We note here that tracking distances between successive iterates is common in the analysis of fixed-point iterations; see, for example, [BB96, CSV14, SS17, BCC21, CC23]. This simple observation combined with the definition of the iteration (FixHal) can then be used to argue about convergence to an approximate fixed point error dependent on $\lambda,$ as follows.
Lemma 1 (Convergence of the Fixed-Stepsize Iteration).
Given ${\bm{x}}_{0}\in{\mathcal{E}}$ and iterates ${\bm{x}}_{k+1}$ defined by (FixHal) for a $\gamma$-Lipschitz operator ${\bm{T}}$ with $\gamma>0,$ we have that for all $k\geq 1,$
$\|{\bm{T}}({\bm{x}}_{k})-{\bm{x}}_{k}\|\leq$ $(1-\lambda)^{k}\gamma^{k}\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|+\frac{\lambda}{1-\lambda}\|{\bm{x}}_{0}-{\bm{x}}_{k}\|.$
If $\gamma\in(0,1]$ and ${\bm{x}}_{*}$ is a fixed point of ${\bm{T}},$ then we further have
$\|{\bm{T}}({\bm{x}}_{k})-{\bm{x}}_{k}\|\leq(1-\lambda)^{k}\gamma^{k}\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|+\frac{2\lambda}{1-\lambda}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|.$ (3)
As a consequence, for $\gamma\in(0,1]$ and given any $\epsilon>0,$ if $\lambda=\frac{\epsilon}{4D_{*}+\epsilon}$, where $D_{*}\geq\|{\bm{x}}_{0}-{\bm{x}}_{*}\|,$ then $\|{\bm{T}}({\bm{x}}_{k})-{\bm{x}}_{k}\|\leq\epsilon$ after at most $k$ iterations, where
$k=\Big{\lceil}\frac{\ln(2\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|/\epsilon)}{\ln(1/(1-\lambda))+\ln(1/\gamma)}\Big{\rceil}={\mathcal{O}}\Big{(}\frac{\ln(\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|/\epsilon)}{\epsilon/D_{*}+\ln(1/\gamma)}\Big{)}.$
Proof.
Observe first that by (FixHal), we have ${\bm{x}}_{1}-{\bm{x}}_{0}=(1-\lambda)({\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}).$ Further, by (1) and $\gamma$-Lipschitzness of ${\bm{T}},$
$\|{\bm{x}}_{k+1}-{\bm{x}}_{k}\|\leq(1-\lambda)\gamma\|{\bm{x}}_{k}-{\bm{x}}_{k-1}\|$ (4)
for any $k\geq 1.$ Applying (4) recursively we thus get
$\|{\bm{x}}_{k+1}-{\bm{x}}_{k}\|\leq(1-\lambda)^{k}\gamma^{k}\|{\bm{x}}_{1}-{\bm{x}}_{0}\|=(1-\lambda)^{k+1}\gamma^{k}\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|.$ (5)
On the other hand, subtracting ${\bm{x}}_{k}$ from (FixHal) gives ${\bm{x}}_{k+1}-{\bm{x}}_{k}=\lambda({\bm{x}}_{0}-{\bm{x}}_{k})+(1-\lambda)({\bm{T}}({\bm{x}}_{k})-{\bm{x}}_{k}),$ or, equivalently after a rearrangement, ${\bm{T}}({\bm{x}}_{k})-{\bm{x}}_{k}=\frac{1}{1-\lambda}({\bm{x}}_{k+1}-{\bm{x}}_{k})-\frac{\lambda}{1-\lambda}({\bm{x}}_{0}-{\bm{x}}_{k}).$ Thus, by the triangle inequality,
$\|{\bm{T}}({\bm{x}}_{k})-{\bm{x}}_{k}\|$ $\leq\frac{1}{1-\lambda}\|{\bm{x}}_{k+1}-{\bm{x}}_{k}\|+\frac{\lambda}{1-\lambda}\|{\bm{x}}_{0}-{\bm{x}}_{k}\|$ (6)
$\leq(1-\lambda)^{k}\gamma^{k}\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|+\frac{\lambda}{1-\lambda}\|{\bm{x}}_{0}-{\bm{x}}_{k}\|,$
where the second inequality is by (5). This gives the first inequality from the lemma statement.
For the second inequality in the lemma statement, we observe by the triangle inequality that $\|{\bm{x}}_{0}-{\bm{x}}_{k}\|\leq\|{\bm{x}}_{0}-{\bm{x}}_{*}\|+\|{\bm{x}}_{k}-{\bm{x}}_{*}\|,$ so it suffices to show that $\|{\bm{x}}_{k}-{\bm{x}}_{*}\|\leq\|{\bm{x}}_{0}-{\bm{x}}_{*}\|,$ for $k\geq 0.$ This follows by induction on $k,$ since using that $\gamma\leq 1$ and ${\bm{T}}({\bm{x}}_{*})={\bm{x}}_{*},$ we have for all $k\geq 0,$
$\|{\bm{x}}_{k+1}-{\bm{x}}_{*}\|$ $\leq\lambda\|{\bm{x}}_{0}-{\bm{x}}_{*}\|+(1-\lambda)\|{\bm{T}}({\bm{x}}_{k})-{\bm{T}}({\bm{x}}_{*})\|$
$\leq\lambda\|{\bm{x}}_{0}-{\bm{x}}_{*}\|+(1-\lambda)\|{\bm{x}}_{k}-{\bm{x}}_{*}\|.$
To bound the number of iterations until $\|{\bm{T}}({\bm{x}}_{k})-{\bm{x}}_{k}\|\leq\epsilon,$ observe first that when $\lambda=\frac{\epsilon}{4D_{*}+\epsilon},$ we have $\frac{2\lambda}{1-\lambda}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|\leq\frac{\epsilon}{2},$ so it suffices to bound the number of iterations $k$ until
$(1-\lambda)^{k}\gamma^{k}\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|\leq\frac{\epsilon}{2}.$
Solving the last inequality for $k,$ we get that any $k\geq\frac{\ln(2\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|/\epsilon)}{\ln(1/(1-\lambda))+\ln(1/\gamma)}$ suffices. Rounding the right-hand side to the larger integer number leads to the bound claimed in the lemma statement. ∎
2.2 Convergence for Mildly Expansive Operators
We can observe from the analysis from the last subsection that for (FixHal) to converge to an approximate fixed point, ${\bm{T}}$ is not required to be non-expansive. In particular, we can allow $\gamma$ to be slightly larger than one, and we informally refer to such operators as being “mildly expansive.” All that is needed to ensure convergence is that (1) the iterates remain in a bounded set so that $\|{\bm{x}}_{0}-{\bm{x}}_{k}\|$ is bounded by some $D<\infty$ and (2) for the step size $\lambda$, we have $(1-\lambda)\gamma<1.$ To keep the definitions and the main message reasonably simple, throughout this subsection and whenever we work with possibly expansive operators, we assume that ${\bm{T}}$ maps a convex compact set of diameter $D<\infty$ to itself (1). In infinite-dimensional spaces, this assumption is necessary to ensure that minimum possible fixed-point error is finitely bounded [Goe73].
Assumption 1.
There exists a compact convex set ${\mathcal{C}}\subset{\mathcal{E}}$ such that ${\bm{T}}:{\mathcal{C}}\to{\mathcal{C}}$ and $\sup_{{\bm{x}},{\bm{y}}\in{\mathcal{C}}}\|{\bm{x}}-{\bm{y}}\|\leq D<\infty.$
Based on this assumption, we further assume that (FixHal) is initialized at ${\bm{x}}_{0}\in{\mathcal{C}}$, so that all iterates remain in ${\mathcal{C}}.$ We now formally prove the claimed result for a mildly expansive (a.k.a. almost nonexpansive [LTT18]) operator ${\bm{T}}.$
Lemma 2 (Convergence of (FixHal) for a mildly expansive operator.).
Given a $\gamma$-Lipschitz operator ${\bm{T}}$ that satisfies 1, $\epsilon>0$, and $\beta\in(0,1),$ consider $\{{\bm{x}}_{k}\}_{k\geq 0}$ initialized at some ${\bm{x}}_{0}\in{\mathcal{C}}$ and updated according to (FixHal) for $k\geq 0$, where $\lambda=\frac{\beta\epsilon/D}{1+\beta\epsilon/D}$ and $D$ is the diameter of ${\mathcal{C}}.$ If $0\leq\gamma<1+\frac{\beta\epsilon}{D},$ then $\|{\bm{T}}({\bm{x}}_{k})-{\bm{x}}_{k}\|\leq\epsilon$ after at most $k=\big{\lceil}\frac{\ln(\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|/((1-\beta)\epsilon))}{-\ln((1-\lambda)\gamma)}\big{\rceil}$ iterations.
Proof.
Applying Lemma 1, we have that for all $k\geq 0,$
$\|{\bm{T}}({\bm{x}}_{k})-{\bm{x}}_{k}\|$ $\leq(1-\lambda)^{k}\gamma^{k}\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|+\frac{\lambda}{1-\lambda}\|{\bm{x}}_{0}-{\bm{x}}_{k+1}\|$
$\leq(1-\lambda)^{k}\gamma^{k}\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|+\beta\epsilon.$
The bound on $\gamma$ from the lemma statement ensures that $(1-\lambda)\gamma<1.$ Thus, to obtain the claimed bound on the number of iterations, we need to determine $k$ that ensures $(1-\lambda)^{k}\gamma^{k}\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|\leq(1-\beta)\epsilon.$ Solving this inequality for $k,$ we get that the inequality is satisfied for $k\geq\frac{\ln(\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|/((1-\beta)\epsilon))}{-\ln((1-\lambda)\gamma)}$, and thus $k=\big{\lceil}\frac{\ln(\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|/((1-\beta)\epsilon))}{-\ln((1-\lambda)\gamma)}\big{\rceil}$ iterations suffice. ∎
Another way of phrasing Lemma 2 is in terms of the attainable fixed-point error for an operator ${\bm{T}}$ with an arbitrary Lipschitz constant $\gamma>1.$ We state this result as the corollary below, which will be particularly useful for discussing this result in infinite-dimensional spaces.
Corollary 1.
Given a $\gamma$-Lipschitz operator ${\bm{T}}$ that satisfies 1, where $\gamma>1$, and $\beta\in(0,1),$ consider $\{{\bm{x}}_{k}\}_{k\geq 0}$ initialized at some ${\bm{x}}_{0}\in{\mathcal{C}}$ and updated according to (FixHal) for $k\geq 0$, where $\lambda=1-\frac{\beta}{\gamma}$. Then, for any $\epsilon\in(0,\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|],$ we have $\|{\bm{T}}({\bm{x}}_{k})-{\bm{x}}_{k}\|\leq\big{(}\frac{\gamma}{\beta}-1\big{)}D+\epsilon$ after at most $k=\big{\lceil}\frac{\ln(\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|/\epsilon)}{\ln(1/\beta)}\big{\rceil}$ iterations. In particular, for $\beta=1-\frac{\epsilon}{2\gamma D}$, we have $\|{\bm{T}}({\bm{x}}_{k})-{\bm{x}}_{k}\|\leq(\gamma-1)D+\epsilon$ after $k={\mathcal{O}}\big{(}\frac{\gamma D\ln(\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|/\epsilon)}{\epsilon}\big{)}$ iterations.
Proof.
Observe first that $(1-\lambda)\gamma=\beta\in(0,1)$ and $\frac{\lambda}{1-\lambda}\|{\bm{x}}_{0}-{\bm{x}}_{k+1}\|\leq\big{(}\frac{\gamma}{\beta}-1\big{)}D.$ Thus, Lemma 1 implies that for all $k\geq 0,$ $\|{\bm{T}}({\bm{x}}_{k})-{\bm{x}}_{k}\|\leq\beta^{k}\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|+\big{(}\frac{\gamma}{\beta}-1\big{)}D.$ Now, for any $\epsilon>0,$ we have that $\beta^{k}\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|\leq\epsilon$ after $k\geq\log_{1/\beta}\big{(}\frac{\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|}{\epsilon}\big{)}=\frac{\ln({\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|}/{\epsilon})}{\ln(1/\beta)}$ iterations. The last bound follows by a direct calculation, choosing $k$ slightly larger so that $\beta^{k}\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|\leq\epsilon/2.$
We now provide a discussion of the obtained results. The entire discussion applies under 1.
Finite-dimensional vector space.
Consider first the case of a finite-dimensional vector space, for instance, $\mathbb{R}^{d}.$ In this case the existence of a fixed point is guaranteed via Brouwer’s fixed point theorem, as discussed before. To approximate the fixed point error to some $\epsilon>0,$ Lemma 2 requires the operator ${\bm{T}}$ to have only very mild expansion, with Lipschitz constant allowed to be only slightly larger than one. This may not seem impressive on its own, since it only allows $\gamma$ to take values up to $1+\frac{\beta\epsilon}{D}$ for $\beta\in(0,1).$ However, it turns out that (up to absolute constants), this is the best we can hope to get with an algorithm that makes a subexponential (in the dimension) number of queries to ${\bm{T}}$, due to the lower bound from [HPV89]. In particular, let $M$ be the Lipschitz constant of ${\bm{T}}-{\bm{\mathrm{Id}}},$ and observe that $\gamma\leq 1+M$ (so $M\geq\gamma-1$) by the triangle inequality. Theorem 2 in [HPV89] implies that for any $\epsilon\in(0,1/10),$ any oracle-based algorithm applied to an operator ${\bm{T}}$ that maps the unit hypercube $[0,1]^{d}$ (so $D=1$), where $d\geq 3$, to itself and is $(1+M)$-Lipschitz continuous with respect to the $\ell_{\infty}$ norm, in the worst case must make at least
$\Big{(}c\Big{(}\frac{1}{\epsilon}-10\Big{)}M\Big{)}^{d-2}$ (7)
queries to ${\bm{T}}$ to reach a vector $\hat{\bm{x}}$ in the unit hypercube such that its fixed point error is at most $\epsilon$ (i.e., $\|{\bm{T}}(\hat{\bm{x}})-\hat{\bm{x}}\|_{\infty}\leq\epsilon$), where $c\geq 10^{-5}$. Here, the important thing about $c$ is that it is an absolute constant, independent of the problem parameters. The authors conjecture that $c$ can in fact be made close to one. Observe that Lemma 2 gives a polynomial bound on the number of queries of (FixHal) for problems described in this lower bound whenever $\gamma\leq 1+{\beta\epsilon}$ (as $D=1$) and $\beta$ is smaller than one by some absolute constant (e.g., we can take $\beta=0.99$ and bound the number of queries by ${\mathcal{O}}\big{(}\frac{\ln(1/\epsilon)}{\epsilon}\big{)}$). On the other hand, the lower bound stated in (7) implies that $M$ must be bounded by an absolute constant times $\epsilon$ for the number of worst-case oracle queries of any algorithm to be subexponential, meaning that there exists an absolute constant $c^{\prime}>0$ such that any algorithm that finds an $\epsilon$-fixed point of a $(1+c^{\prime}\epsilon)$-Lipschitz operator must make an exponential in the dimension number of queries in the worst case. Thus, the result from Lemma 2 cannot be improved, modulo the constant multiplying $\epsilon$ in the bound on $\gamma.$
While the above discussion considered the case $D=1,$ we note that, by a simple rescaling argument, the dependence on a more general value of $D>0$ as in our bound is similarly unavoidable (up to absolute constants, as discussed above). To see this, consider the following. Let ${\bm{F}}={\bm{T}}-{\bm{\mathrm{Id}}}$, so that $M$ is the Lipschitz constant of ${\bm{F}},$ for ${\bm{F}}$ mapping the unit hypercube in ${\mathbb{R}}^{d}$ to itself, as in the above discussion. Define $\widehat{\bm{F}}$—an operator mapping the hypercube $[0,D]^{d}$ to itself—via $\widehat{\bm{F}}({\bm{x}}):={\bm{F}}({\bm{x}}/D)$. It is immediate that the Lipschitz constant of $\widehat{\bm{F}}$ is $\widehat{M}=M/D$ and that for any $\epsilon>0$ and any ${\bm{x}}$ from the side-$D$ hypercube we have $\|\widehat{\bm{F}}({\bm{x}})\|_{\infty}\leq\epsilon$ if and only if $\|{\bm{F}}({\bm{x}}/D)\|_{\infty}\leq\epsilon$. Now, the lower bound stated in (7) implies that for $d\geq 3$ and any $\epsilon\in(0,1/10)$, any algorithm based on oracle queries to $\widehat{\bm{F}}$ must make at least
$\Big{(}c\Big{(}\frac{1}{\epsilon}-10\Big{)}\widehat{M}D\Big{)}^{d-2}$ (8)
queries to $\widehat{\bm{F}}$ in the worst case to output a point ${\bm{x}}$ with error $\|\widehat{\bm{F}}({\bm{x}})\|_{\infty}\leq\epsilon$. A consequence of this statement is that there exists an absolute constant $c^{\prime}$ such that no (oracle query-based) algorithm can output a point with fixed-point error $\epsilon\in(0,1/10)$ for an arbitrary $(1+c^{\prime}\epsilon/D)$-Lipschitz operator mapping a compact convex set of diameter $D$ to itself without incurring an exponential dependence in the dimension (assuming $d\geq 3$) for the number of required oracle queries to the operator. Thus, for $d\geq 3,$ the result from Lemma 2 cannot be improved (modulo the absolute constant factor multiplying $\epsilon/D$) without either making additional assumptions about the operator ${\bm{T}}$ or incurring exponential complexity in the dimension. Put differently, for $\gamma>1$ by any however small constant positive amount, no oracle-based algorithm can guarantee convergence to an arbitrarily small fixed-point error in polynomial time.
Infinite-dimensional space.
For infinite-dimensional Banach spaces $({\mathcal{E}},\left\lVert\cdot\right\rVert)$, the existence of a fixed point is not guaranteed under 1, even if the operator ${\bm{T}}$ is nonexpansive [Bro65, Göh65, Kir65]. Instead, for $\gamma$-Lipschitz operators ${\bm{T}}$ with $\gamma>1,$ it is only possible to guarantee that $\inf_{{\bm{x}}\in{\mathcal{C}}}\|{\bm{T}}({\bm{x}})-{\bm{x}}\|\leq(1-1/\gamma)D,$ and this inequality is tight in general [Goe73]. [^7] Since the problem is no longer guaranteed to have a fixed point, the problem now becomes to find solutions with minimum value of $\|{\bm{T}}({\bm{x}})-{\bm{x}}\|,$ which is also referred to as the “minimal displacement problem;” see [Goe73]. In other words, the best fixed-point error we can hope to achieve with any algorithm would be of the form $(1-1/\gamma)D+\epsilon,$ for $\epsilon>0.$ Based on Corollary 1, (FixHal) can attain fixed-point error $(\gamma-1)D+\epsilon$ after $k={\mathcal{O}}\big{(}\frac{\gamma\ln(\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|/\epsilon)}{\epsilon}\big{)}$ iterations, which is off by a multiplicative factor $\gamma$ and additive error $\epsilon$ compared to the bound on the minimal error $(1-1/\gamma)D,$ since $(\gamma-1)D+\epsilon=\gamma(1-1/\gamma)D+\epsilon$. To the best of our knowledge, this is the first result that gives any nontrivial bound on the fixed-point error of general Lipschitz operators using a finite number of operator evaluations, despite a relevant open question being raised in [BB96].
We note here that for sufficiently large $\gamma,$ the lower bound from [HPV89] stated in (8) for the hypercube of side $D$ in ${\mathbb{R}}^{d}$ ($d\geq 3$) does not preclude results of the form $(1-1/\gamma)D+\epsilon,$ for $\epsilon>0,$ on the basis of such problems being included in the more general fixed-point problems in Banach spaces $({\mathcal{E}},\left\lVert\cdot\right\rVert)$. The reason is that the shift by 10 in (8) makes the lower bound meaningful only for sufficiently small values of the target fixed-point error. It is thus an interesting question whether error closer to $(1-1/\gamma)D+\epsilon$ than what we obtained is algorithmically attainable in general Banach spaces. This was conjectured true for the KM iteration in [BB96]; however, to the best of our knowledge, such a result has not been established to date.
2.3 Extension to Geodesic Metric Spaces
Even though our results are presented for normed vector spaces, they can be generalized to certain geodesic metric spaces. The reason is that very few properties of the norm are used in the analysis of (FixHal) and algorithms developed based on it in the upcoming section; namely, in addition to properties defining any metric, only convexity of the norm (or a metric) is needed. This means that the results extend to positively curved geodesic metric spaces known as Busemann spaces, which are fully characterized by convex metrics on every geodesic (see the textbook [Bač14] for relevant definitions and, in particular, Proposition 1.1.5 for the characterization of Busemann geodesic metric spaces via convexity of the metric).
For concreteness, let $({\mathcal{G}},d)$ denote a geodesic metric space. With a slight abuse of notation confined to this subsection, we use $d$ to denote the metric associated with the space $({\mathcal{G}},d)$. Let ${\bm{T}}:{\mathcal{G}}\to{\mathcal{G}}$ be an operator that is $\gamma$-Lipschitz for some $\gamma>0,$ which for the present setting means that $d({\bm{T}}({\bm{x}}),{\bm{T}}({\bm{y}}))\leq\gamma\,d({\bm{x}},{\bm{y}})$ for any ${\bm{x}},{\bm{y}}\in{\mathcal{G}}.$
For iterates of (FixHal) to remain in ${\mathcal{G}},$ the iteration update for a step size $\lambda\in(0,1)$ is written as
${\bm{x}}_{k+1}=\lambda{\bm{x}}_{0}\oplus(1-\lambda){\bm{T}}({\bm{x}}_{k}),$ (9)
where the above equation is interpreted as ${\bm{x}}_{k+1}$ being the element of ${\mathcal{G}}$ on a geodesic from ${\bm{x}}_{0}$ to ${\bm{T}}({\bm{x}}_{k})$ such that $d({\bm{x}}_{k+1},{\bm{x}}_{0})=(1-\lambda)d({\bm{T}}({\bm{x}}_{k}),{\bm{x}}_{0}).$
For concreteness, we show how to obtain a counterpart to Lemma 1 for a $\gamma$-Lipschitz operator ${\bm{T}}$ on a Busemann space $({\mathcal{G}},d).$ With this argument, all other results in this paper can be generalized to Busemann spaces in a straightforward manner, which is omitted for brevity.
Lemma 3 (Convergence of the Fixed-Stepsize Iteration in a Busemann Space).
Let $({\mathcal{G}},d)$ be a Busemann space and ${\bm{T}}:{\mathcal{G}}\to{\mathcal{G}}$ be an operator that is $\gamma$-Lipschitz on ${\mathcal{G}}$ for some $\gamma>0.$ Given ${\bm{x}}_{0}\in{\mathcal{G}}$ and iterates ${\bm{x}}_{k+1}$ defined by (9) for $k\geq 0$, we have that for all $k\geq 1,$
$d({\bm{T}}({\bm{x}}_{k}),{\bm{x}}_{k})\leq$ $(1-\lambda)^{k}\gamma^{k}d({\bm{T}}({\bm{x}}_{0}),{\bm{x}}_{0})+\frac{\lambda}{1-\lambda}d({\bm{x}}_{0},{\bm{x}}_{k}).$
If $\gamma\in(0,1]$ and ${\bm{x}}_{*}$ is a fixed point of ${\bm{T}},$ then we further have
$d({\bm{T}}({\bm{x}}_{k}),{\bm{x}}_{k})\leq(1-\lambda)^{k}\gamma^{k}d({\bm{T}}({\bm{x}}_{0}),{\bm{x}}_{0})+2\frac{\lambda}{1-\lambda}d({\bm{x}}_{0},{\bm{x}}_{*}).$ (10)
As a consequence for $\gamma\in(0,1]$, given any $\epsilon>0,$ if $\lambda=\frac{\epsilon}{4D_{*}+\epsilon}$, where $D_{*}\geq d({\bm{x}}_{0},{\bm{x}}_{*}),$ then $d({\bm{T}}({\bm{x}}_{k}),{\bm{x}}_{k})\leq\epsilon$ after at most $k$ iterations, where
$k=\Big{\lceil}\frac{\ln(2d({\bm{T}}({\bm{x}}_{0}),{\bm{x}}_{0})/\epsilon)}{\ln(1/(1-\lambda))+\ln(1/\gamma)}\Big{\rceil}={\mathcal{O}}\Big{(}\frac{\ln(d({\bm{T}}({\bm{x}}_{0}),{\bm{x}}_{0})/\epsilon)}{\epsilon/D_{*}+\ln(1/\gamma)}\Big{)}.$
Proof.
Similar to the proof of Lemma 1, we first bound the distance between successive iterates. For $k=0$, we have, by the definition of the iteration (9), that $d({\bm{x}}_{1},{\bm{x}}_{0})=(1-\lambda)d({\bm{T}}({\bm{x}}_{0}),{\bm{x}}_{0})$. For $k\geq 1$, by convexity of the metric $d$ and by (9),
$d({\bm{x}}_{k+1},{\bm{x}}_{k})\leq\;$ $\lambda d({\bm{x}}_{0},{\bm{x}}_{0})+(1-\lambda)d({\bm{T}}({\bm{x}}_{k}),{\bm{T}}({\bm{x}}_{k-1}))$
[…]
where the last inequality is by $d({\bm{x}},{\bm{x}})=0$ for any ${\bm{x}}\in{\mathcal{G}}$ (by the definition of a metric) and by $d({\bm{T}}({\bm{x}}_{k}),{\bm{T}}({\bm{x}}_{k-1}))\leq\gamma d({\bm{x}}_{k},{\bm{x}}_{k-1})$, which holds by the definition of $\gamma$-Lipschitzness of ${\bm{T}}$.
Applying the last inequality recursively and using that $d({\bm{x}}_{1},{\bm{x}}_{0})=(1-\lambda)d({\bm{T}}({\bm{x}}_{0}),{\bm{x}}_{0})$, we thus have
$d({\bm{x}}_{k+1},{\bm{x}}_{k})\leq(1-\lambda)^{k+1}\gamma^{k}d({\bm{T}}({\bm{x}}_{0}),{\bm{x}}_{0}).$ (11)
Now once again using the definition of the iteration (9) and convexity of $d$, we get
$d({\bm{T}}({\bm{x}}_{k+1}),{\bm{x}}_{k+1})\leq\;$ $\lambda d({\bm{x}}_{0},{\bm{T}}({\bm{x}}_{k+1}))+(1-\lambda)d({\bm{T}}({\bm{x}}_{k+1}),{\bm{T}}({\bm{x}}_{k}))$
[…]
where the second inequality uses $\gamma$-Lipschitzness of ${\bm{T}}$, the third inequality uses (11), and the last inequality is by applying the triangle inequality as $d({\bm{x}}_{0},{\bm{T}}({\bm{x}}_{k+1}))\leq d({\bm{x}}_{0},{\bm{x}}_{k+1})+d({\bm{T}}({\bm{x}}_{k+1}),{\bm{x}}_{k+1})$. Grouping the like terms in (12) and simplifying now gives the first claim.
The rest of the lemma is proved in the same way as Lemma 1, as the only properties of the norm used there are the triangle inequality (which holds for an arbitrary metric by definition) and convexity of the metric (which holds for metrics in Busemann spaces). Thus, we omit it for brevity. ∎
3 Gradual Halpern Algorithm and its Adaptive Version
Although it is possible to use the fixed-step Halpern iteration (FixHal) to approximately solve fixed-point equations with contractive, nonexpansive, or even mildly expansive operators, there are multiple downsides to its direct use. First, even if we knew everything about the problem, we fixed the target approximation $\epsilon>0$, and set the step size $\lambda=\lambda(\epsilon)$ accordingly, the iteration complexity obtained from Lemma 1 would be suboptimal by a factor $\ln\big{(}\frac{\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|}{\epsilon})$ for nonexpansive operators [DW22, PR22]. Second, it is generally undesirable to set the target error in advance as done in this case, because, in the worst case, the algorithm run for no matter how long would never converge to a fixed point error lower than some positive threshold dependent on $\epsilon.$ Third, the value of $\lambda$ prescribed by Lemma 1 requires an upper bound on the iterate distance $\|{\bm{x}}_{k+1}-{\bm{x}}_{0}\|\leq D$ to be useful, which is a strong assumption particularly when the operator ${\bm{T}}$ is nonexpansive (i.e., when $\gamma\leq 1$). Finally, since the algorithm can provide guarantees for any value of $\gamma\in(0,1+\epsilon/D)$, it would be useful to have an algorithm that can automatically adapt to the value of $\gamma,$ possibly even locally, without explicit prior knowledge about local or global values of $\gamma.$ In this section, we provide and discuss algorithms that address these issues.
For clarity of exposition, we handle separately the cases in which the operator ${\bm{T}}$ is nonexpansive and in which it is mildly expansive, and build from the setting where we know the diameter of the set containing algorithm iterates to a fully parameter-free setting.
3.1 Gradual Halpern Algorithm (GHAL)
We consider first a simple setting in which we have an upper estimate $D$ for the diameter of the set containing algorithm’s iterates. In particular, for nonexpansive operators ($\gamma\leq 1$), we have $D\leq 2\|{\bm{x}}_{0}-{\bm{x}}_{*}\|$, while under 1, $D$ is the diameter of the set ${\mathcal{C}}$ for which ${\bm{T}}:{\mathcal{C}}\to{\mathcal{C}}.$ The pseudocode for the algorithm (Gradual Halpern Algorithm—GHAL) is provided in Algorithm 1. In the algorithm, there is no apparent benefit to choosing $\beta$ to be anything other than an absolute constant, so in the analysis below it helps to regard $\beta$ as some number between zero and one, such as, for example, $0.1,0.5,$ or $0.9$
Algorithm 1 Gradual Halpern Algorithm (GHAL)
1:Input: $\epsilon>0,{\bm{x}}_{0},D<\infty,$ $\beta\in(0,1),\beta^{\prime}\in(0,1)$
2:Initialization: $\epsilon_{0}\leftarrow\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|,\,k\leftarrow 0,\,\hat{\bm{x}}_{0}\leftarrow{\bm{x}}_{0}$
3:while $\|{\bm{T}}(\hat{\bm{x}}_{k})-\hat{\bm{x}}_{k}\|>\epsilon$ do
4:     $k\leftarrow k+1$
5:     $\epsilon_{k}\leftarrow\beta\epsilon_{k-1}$
6:     $\lambda_{k}\leftarrow\frac{\beta\epsilon_{k}/D}{1+\beta\epsilon_{k}/D}$
7:     ${\bm{y}}_{0}=\hat{\bm{x}}_{k-1}$, $j=0$triangleright Initialize (FixHal)
8:     while $\|{\bm{T}}({\bm{y}}_{j})-{\bm{y}}_{j}\|>\epsilon_{k}$ do $\triangleright$ Run (FixHal) until target error is reached
9:         ${\bm{y}}_{j+1}\leftarrow\lambda_{k}{\bm{y}}_{0}+(1-\lambda_{k}){\bm{T}}({\bm{y}}_{j})$
10:         if $j\geq 2$ and $\|{\bm{y}}_{j+1}-{\bm{y}}_{j}\|\geq(1-\beta^{\prime}\lambda_{k})\|{\bm{y}}_{j}-{\bm{y}}_{j-1}\|$ then $\triangleright$ Safeguard for potentially expansive operators
11:              Halt the algorithm and return $\hat{\bm{x}}_{k}=\operatorname*{arg\,min}\{\|{\bm{T}}({\bm{y}}_{j})-{\bm{y}}_{j}\|,\|{\bm{T}}({\bm{y}}_{0})-{\bm{y}}_{0}\|\}$
12:         end if
13:         $j\leftarrow j+1$
14:     end while
15:     $\hat{\bm{x}}_{k}={\bm{y}}_{j}$
16:end while
17:return $\hat{\bm{x}}_{k}$
Proposition 1.
Let ${\bm{T}}$ be a $\gamma$-Lipschitz operator and let $\hat{\bm{x}}_{k}$ be the output of Algorithm 1 initialized at some ${\bm{x}}_{0}$ and run for $\beta\in(0,1)$ such that $1/\beta,1/(1-\beta)$ are absolute constants independent of the problem parameters. Suppose $\left\lVert{\bm{y}}_{j}-{\bm{y}}_{0}\right\rVert\leq D$ for any iterates ${\bm{y}}_{j}$ within Algorithm 1. If $\gamma\in(0,1],$ then $\|{\bm{T}}(\hat{\bm{x}}_{k})-\hat{\bm{x}}_{k}\|\leq\epsilon$ and the algorithm makes
${\mathcal{O}}\Big{(}\min\Big{\{}{\ln(\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|/\epsilon)}\frac{1+\ln(1/\gamma)}{\ln(1/\gamma)},\,\ln(1/\gamma)+\frac{\ln(\gamma\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|/\epsilon)}{\ln(1/\gamma)},\,\frac{D}{\epsilon}\big{\}}\big{)}$
calls to (FixHal) iterations (and consequently the same number of oracle calls to ${\bm{T}}$).
If $\gamma\in(1,1+\frac{\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|}{D}\frac{\beta^{2}}{1-\beta^{\prime}}],$ then Algorithm 1 outputs a point $\hat{\bm{x}}_{k}$ with $\|{\bm{T}}(\hat{\bm{x}}_{k})-\hat{\bm{x}}_{k}\|\leq\max\{\epsilon,\bar{\epsilon}\}$, where $\bar{\epsilon}=\frac{D}{\beta^{2}}\frac{\gamma-1}{1-\beta^{\prime}}$, after at most ${\mathcal{O}}(\min\{\frac{D}{\epsilon\beta^{\prime}},\,\frac{\beta(1-\beta^{\prime})}{\gamma-1}\})$ oracle calls to ${\bm{T}}.$
Proof.
We carry out the proof by separately handling the cases in which the operator is nonexpansive and in which it is mildly expansive.
Case 1: Nonexpansive operator. We first consider the case in which $\gamma\in(0,1]$ and observe that it is impossible for the algorithm to halt due to the condition in Line 8, since we established in the proof of Lemma 1 that for nonexpansive operators, (FixHal) must be contracting distances between successive iterates by a factor at most $1-\lambda_{k}$. The claim $\|{\bm{T}}(\hat{\bm{x}}_{k})-\hat{\bm{x}}_{k}\|\leq\epsilon$ thus follows by the outer while loop condition—for the while loop to terminate and the algorithm to halt, this condition must hold. Because $\epsilon_{k}$ reduces by a factor $\beta\in(0,1)$ in each sequence of calls to (FixHal) within the inner while loop, there are at most $\lceil\log_{1/\beta}(\epsilon_{0}/(\beta\epsilon))\rceil$ iterations of the while loop. Using Lemma 1, each iteration $k$ of the while loop makes at most $\Big{\lceil}\frac{\log_{1/\beta}(\|{\bm{T}}(\hat{\bm{x}}_{k-1})-\hat{\bm{x}}_{k-1}\|/(\beta\epsilon_{k}))}{\log_{1/\beta}(1/(1-\lambda_{k}))+\log_{1/\beta}(1/\gamma)}\Big{\rceil}={\mathcal{O}}(1+\frac{1}{\epsilon_{k}/D+\ln(1/\gamma)})={\mathcal{O}}(1+\frac{1}{\epsilon_{0}\beta^{k}/D+\ln(1/\gamma)})$ calls to (FixHal), where we used that $\|{\bm{T}}(\hat{\bm{x}}_{k-1})-\hat{\bm{x}}_{k-1}\|\leq\epsilon_{k-1}=\epsilon_{k}/\beta$ and that $1/\beta$ is an absolute constant. Hence the total number of calls to (FixHal), each requiring a single oracle call to ${\bm{T}}$, is
${\mathcal{O}}\bigg{(}\sum_{k=1}^{\log_{1/\beta}(\epsilon_{0}/\epsilon)}\Big{(}1+\frac{1}{\epsilon_{0}\beta^{k}/D+\ln(1/\gamma)}\Big{)}\bigg{)}$ $={\mathcal{O}}\bigg{(}\min\Big{\{}\ln(\epsilon_{0}/\epsilon)\frac{1+\ln(1/\gamma)}{\ln(1/\gamma)},\,\frac{D}{\epsilon_{0}}\Big{(}\frac{1}{\beta}\Big{)}^{\log_{1/\beta}(\epsilon_{0}/\epsilon)}\Big{\}}\bigg{)}$
$={\mathcal{O}}\bigg{(}\min\Big{\{}\ln(\epsilon_{0}/\epsilon)\frac{1+\ln(1/\gamma)}{\ln(1/\gamma)},\,\frac{D}{\epsilon}\Big{\}}\bigg{)}.$
Now consider $\gamma<\beta/2.$ Observe that, by the above argument, no iteration of the while loop can result in more than a constant number of iterations, since ${\mathcal{O}}(1+1/\ln(1/\gamma))$ is bounded by a constant in this case. We claim that in this case, after an initial small number of while loop iterations, the inner while loop of Algorithm 1 makes only one call to (FixHal), and, moreover, the contraction of the fixed point error quickly reaches the order of $\gamma$. Let ${\bm{y}}_{0}=\hat{\bm{x}}_{k-1},$ ${\bm{y}}_{1}=\lambda_{k}{\bm{y}}_{0}+(1-\lambda_{k}){\bm{T}}({\bm{y}}_{0}).$ Then, rearranging the ${\bm{y}}_{1}$ update equation, we get that
${\bm{y}}_{1}-{\bm{y}}_{0}=(1-\lambda_{k})({\bm{T}}({\bm{y}}_{0})-{\bm{y}}_{0})$ $\Rightarrow\|{\bm{y}}_{1}-{\bm{y}}_{0}\|\leq(1-\lambda_{k})\|{\bm{T}}({\bm{y}}_{0})-{\bm{y}}_{0}\|,$ (13)
and, further
$(1-\lambda_{k})({\bm{y}}_{1}-{\bm{T}}({\bm{y}}_{1}))=\lambda_{k}({\bm{y}}_{0}-{\bm{y}}_{1})+(1-\lambda_{k})({\bm{T}}({\bm{y}}_{0})-{\bm{T}}({\bm{y}}_{1})),$
which, using the triangle inequality, $\gamma$-contractivity of ${\bm{T}}$, and (13), implies
$\|{\bm{y}}_{1}-{\bm{T}}({\bm{y}}_{1})\|$ $\leq\Big{(}\frac{\lambda_{k}}{1-\lambda_{k}}+\gamma\Big{)}\|{\bm{y}}_{1}-{\bm{y}}_{0}\|$
$\leq\big{(}\lambda_{k}+(1-\lambda_{k})\gamma\big{)}\|{\bm{T}}({\bm{y}}_{0})-{\bm{y}}_{0}\|.$ (14)
Now, since $\|{\bm{T}}({\bm{y}}_{0})-{\bm{y}}_{0}\|\leq\epsilon_{k-1}$ and calls to (FixHal) in the inner while loop terminate as soon as $\|{\bm{y}}_{j}-{\bm{T}}({\bm{y}}_{j})\|\leq\epsilon_{k}=\beta\epsilon_{k-1},$ a sufficient condition for the while loop to terminate in one iteration is that $\lambda_{k}+(1-\lambda_{k})\gamma\leq\beta.$ Equivalently, this holds if $\lambda_{k}\leq\frac{\beta-\gamma}{1-\gamma}.$ More generally, we have that $\bar{\lambda}_{k}:=\frac{\lambda_{k}}{1-\lambda_{k}}\leq\gamma$ is sufficient to ensure that this while loop terminates in one iteration and $\lambda_{k}+(1-\lambda_{k})\gamma<2\gamma.$ This means that once $\bar{\lambda}_{k}\leq\gamma$, each iteration of Algorithm 1 contracts the fixed point error by a factor at most $2\gamma.$
By the definition of $\lambda_{k},$ we have that $\bar{\lambda}_{k}:=\frac{\lambda_{k}}{1-\lambda_{k}}=\frac{\beta\epsilon_{k}}{D}$, thus the condition $\bar{\lambda}_{k}\leq\gamma$ is equivalent to $\epsilon_{k}\leq D\gamma/\beta.$ The total number of iterations until this value of $\epsilon_{k}$ is reached (of which, recall, each makes a constant number of calls to (FixHal)) is ${\mathcal{O}}(\log_{1/\beta}(\epsilon_{0}/\epsilon_{k}))={\mathcal{O}}\big{(}\ln\big{(}\frac{\epsilon_{0}\beta}{D\gamma}\big{)}\big{)}.$ All subsequent iterations contract the fixed point error by a factor at most $2\gamma,$ so the remaining number of iterations until the algorithm halts is ${\mathcal{O}}\big{(}\frac{\ln(D\gamma/\epsilon)}{\ln(1/\gamma)}\big{)}.$ Observe here that, by the triangle inequality and $\gamma$-contractivity of ${\bm{T}},$ we can conclude that $\epsilon_{0}=\|{\bm{x}}_{0}-{\bm{T}}({\bm{x}}_{0})\|\geq\|{\bm{x}}_{0}-{\bm{x}}_{*}\|-\|{\bm{T}}({\bm{x}}_{0})-{\bm{T}}({\bm{x}}_{*})\|\geq(1-\gamma)\|{\bm{x}}_{0}-{\bm{x}}_{*}\|.$ Thus, if $D={\mathcal{O}}(\|{\bm{x}}_{0}-{\bm{x}}_{*}\|),$ then we also have $D={\mathcal{O}}(\epsilon_{0}),$ and the iteration/oracle complexity in this case simplifies to
${\mathcal{O}}\Big{(}\ln({1}/{\gamma})+\frac{\ln(\gamma\epsilon_{0}/\epsilon)}{\ln(1/\gamma)}\Big{)}.$
To complete the proof of the first case, it remains to recall that $\epsilon_{0}=\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|$ and observe that when $\gamma$ is close to one (but smaller than one), the first two terms in the oracle complexity bound are of the same order.
Case 2: Mildly expansive operator. Suppose now that $\gamma\in(1,2)$. Then it is possible for the algorithm to halt either due to the exit condition of the outer while loop or due to the safeguard condition in Line 8. If the algorithm halts because of the while loop exit condition, then it outputs an $\hat{\bm{x}}_{k}$ with the fixed point error $\|{\bm{T}}(\hat{\bm{x}}_{k})-\hat{\bm{x}}_{k}\|\leq\epsilon.$ Moreover, because in this case the halting condition from Line 8 is never reached, we have that the (FixHal) iterations in the inner loop are always contracting distances by a factor at most $1-\beta^{\prime}\lambda_{k},$ and so we can carry out the same argument as in Case 1 to bound the total number of iterations by ${\mathcal{O}}(\frac{D}{\epsilon\beta^{\prime}}).$
Now suppose that Algorithm 1 halts due to the safeguard condition in Line 8, and suppose that this happens in iteration $k$. The solution output by the algorithm, by the condition in Line 9, has fixed-point error no worse than $\|{\bm{T}}(\hat{\bm{x}}_{k-1})-\hat{\bm{x}}_{k-1}\|\leq\epsilon_{k-1}.$ Because, by (FixHal), we have for $j\geq 2$ that
$\|{\bm{y}}_{j+1}-{\bm{y}}_{j}\|=(1-\lambda_{k})\|{\bm{T}}({\bm{y}}_{j})-{\bm{T}}({\bm{y}}_{j-1})\|\leq(1-\lambda_{k})\gamma\|{\bm{y}}_{j}-{\bm{y}}_{j-1}\|,$
it must be $(1-\lambda_{k})\gamma>1-\beta^{\prime}\lambda_{k},$ or, equivalently, $\lambda_{k}<\frac{\gamma-1}{\gamma-\beta^{\prime}}$. Since $\lambda_{k}=\frac{\beta\epsilon_{k}/D}{1+\beta\epsilon_{k}/D},$ we also have $\epsilon_{k}<\frac{D}{\beta}\frac{\gamma-1}{1-\beta^{\prime}}.$ Thus, the fixed-point error of the solution output by the algorithm is at most $\epsilon_{k-1}=\frac{\epsilon_{k}}{\beta}<\frac{D}{\beta^{2}}\frac{\gamma-1}{1-\beta^{\prime}}\leq\epsilon_{0}$. Since $\epsilon_{k}$ decreases by a factor $\beta$ in each outer while loop iteration, there are at most $\lceil\log_{1/\beta}(\frac{\epsilon_{0}}{\epsilon_{k}})\rceil=\lceil\log_{1/\beta}(\frac{\epsilon_{0}}{D}\frac{\beta(1-\beta^{\prime})}{\gamma-1})\rceil$ such iterations. Since each iteration makes at most ${\mathcal{O}}(\frac{1}{\beta^{\prime}\lambda_{k}})={\mathcal{O}}(\frac{D}{\beta^{\prime}\beta^{k}\epsilon_{0}})$ calls to (FixHal), the total number of calls to (FixHal) (and, consequently, the total number of oracle calls to ${\bm{T}}$) that Algorithm 1 makes is ${\mathcal{O}}(\frac{\beta(1-\beta^{\prime})}{\beta^{\prime}(\gamma-1)})$. ∎
A few remarks are in order here. First, we can observe that Algorithm 1 simultaneously recovers the optimal complexity bounds for nonexpansive and contractive operators in high dimensions, as long as $\gamma$ is not very close to zero, since in that case the complexity would be $\Theta\big{(}\frac{\ln(\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|/\epsilon)}{\ln(1/\gamma)}\big{)}$ [Sik09]. Further, because the step size parameter $\lambda_{k}$ is adjusted gradually, the algorithm can adapt its convergence speed if the operator happens to be locally contractive (but globally nonexpansive). This will also be shown numerically in Section 5.
Second, GHAL is guaranteed to converge to an approximate fixed-point of ${\bm{T}}$ even if ${\bm{T}}$ is not guaranteed to be non-expansive, but only $\gamma$-Lipschitz, with $\gamma\in(1,1+\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|/D),$ where the approximation error scales with the diameter of the set $D$ and with $\gamma-1.$ Note here that the assumption that $\gamma<1+\|{\bm{T}}({\bm{x}}_{0}-{\bm{x}}_{0}\|/D$ is to avoid trivialities. If $\gamma\geq 1+\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|/D$, then the initial solution has error $\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|\leq(\gamma-1)D.$ Since the algorithm cannot output a point with fixed-point error higher than the initial one, the error in this case is guaranteed to be at most $(\gamma-1)D\leq{\overline{\epsilon}}.$ The number of oracle queries is no higher than constant, following the same argument as in the proof of Proposition 1.
Finally, we note that for the purpose of the analysis, the algorithm is stated in a way that guarantees finite halting time when a solution with target error is reached. If the algorithm were to, instead, be called for a fixed number of iterations (as will be done in numerical examples in Section 5), there is no need to halt early. Instead, one can simply replace the “halt” in Line 9 with a “break” for the inner while loop and revert to the previous value of $\lambda_{k-1}$ for the step size. Because Line 9 can only be reached for $\lambda_{k}<\frac{\gamma-1}{\beta^{\prime}-1},$ we can retain the same approximation error guarantee for the algorithm output.
3.2 Adaptive Gradual Halpern Algorithm (AdaGHAL)
A downside of Algorithm 1 is that it is not fully parameter-free since it requires knowledge of $D$ to set the step sizes $\lambda_{k}.$ This assumption can in fact be relaxed by carrying out a more careful analysis. The algorithm is summarized in Algorithm 2. We analyze Algorithm 2 assuming $\gamma\in(0,1];$ however, it seems plausible that the result can be generalized to the case where ${\bm{T}}$ is mildly expansive by adding a safeguard as in Algorithm 1. For clarity of exposition, we omit this generalization and state the algorithm below with $\beta=1/2$ to reduce the notational overhead. We note however, that the same argument and convergence bounds apply if any occurrence of number $2$ in Algorithm 2 is replaced by $1/\beta$ for $\beta\in(0,1)$ and not trivially close to either 0 or 1.
Algorithm 2 Adaptive Gradual Halpern Algorithm (AdaGHAL)
1:Input: $\epsilon>0,{\bm{x}}_{0}$
2:Initialization: $\epsilon_{0}\leftarrow\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|,\,k\leftarrow 0,\,\hat{\bm{x}}_{0}\leftarrow{\bm{x}}_{0}$, $D\leftarrow\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|$
3:while $\|{\bm{T}}(\hat{\bm{x}}_{k})-\hat{\bm{x}}_{k}\|>\epsilon$ do
4:     $k\leftarrow k+1$
5:     $\epsilon_{k}\leftarrow\epsilon_{k-1}/2$
6:     $\lambda_{k}\leftarrow\frac{\epsilon_{k}/(2D)}{1+\epsilon_{k}/(2D)}$
7:     ${\bm{y}}_{0}\leftarrow\hat{\bm{x}}_{k-1}$, $j\leftarrow 0$
8:     while $\|{\bm{T}}({\bm{y}}_{j})-{\bm{y}}_{j}\|>\epsilon_{k}$ do
9:         ${\bm{y}}_{j+1}\leftarrow\lambda_{k}{\bm{y}}_{0}+(1-\lambda_{k}){\bm{T}}({\bm{y}}_{j})$
10:         if $\|{\bm{y}}_{j}-{\bm{y}}_{0}\|\leq D$ then
11:              $j\leftarrow j+1$
12:         else
13:              $D\leftarrow 2D$
14:              $\lambda_{k}\leftarrow\frac{\epsilon_{k}/(2D)}{1+\epsilon_{k}/(2D)}$
15:              ${\bm{y}}_{j}\leftarrow\operatorname*{arg\,min}\big{\{}\|{\bm{T}}({\bm{y}}_{j+1})-{\bm{y}}_{j+1}\|,\|{\bm{T}}({\bm{y}}_{0})-{\bm{y}}_{0}\|\big{\}}$
16:         end if
17:     end while
18:     $\hat{\bm{x}}_{k}\leftarrow{\bm{y}}_{j}$
19:end while
20:return $\hat{\bm{x}}_{k}$
Theorem 1.
Algorithm 2 applied to a $\gamma$-contractive operator ${\bm{T}}$ for $\gamma\in(0,1]$ with error parameter $\epsilon>0$ and initial point ${\bm{x}}_{0}$ outputs a point $\hat{\bm{x}}_{k}$ such that $\|{\bm{T}}(\hat{\bm{x}}_{k})-\hat{\bm{x}}_{k}\|\leq\epsilon.$ The total number of oracle queries to ${\bm{T}}$ that the algorithm makes is ${\mathcal{O}}\big{(}\min\big{\{}\ln(\|{\bm{x}}_{0}-{\bm{x}}_{*}\|/\epsilon)\frac{1+\ln(1/\gamma)}{\ln(1/\gamma)},\,\ln(1/\gamma)+\frac{\ln(\gamma\|{\bm{x}}_{0}-{\bm{x}}_{*}\|/\epsilon)}{\ln(1/\gamma)},\,\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{\epsilon}\big{\}}\big{)}.$
Proof.
The claim that $\|{\bm{T}}(\hat{\bm{x}}_{k})-\hat{\bm{x}}_{k}\|\leq\epsilon$ is immediate, as it corresponds to the exit condition of the outer while loop in Algorithm 2. Thus, as long as the algorithm terminates, the output point $\hat{\bm{x}}_{k}$ must have the target error. In the remainder of the proof, we argue that the algorithm terminates after the claimed number of oracle queries to ${\bm{T}}.$
Let us look at all the iterations that happen between two consecutive updates to the value of $\lambda_{k}.$ The value of $\overline{\lambda}_{k}=\frac{\lambda_{k}}{1-\lambda_{k}}$ always decreases by a factor of 2. This can happen for one of the two possible reasons: (1) because the inner while loop was completed, in which case the upper bound on fixed point error, $\epsilon_{k}$, was halved; or (2) because a point ${\bm{y}}_{j}$ with $\|{\bm{y}}_{j}-{\bm{y}}_{0}\|>D$ was reached, in which case the estimate $D$ gets doubled.
Now consider all calls to ${\bm{T}}$ and let us partition those calls into groups between two consecutive updates to the step size $\lambda_{k}.$ We bound the total number of calls to the oracle by separately bounding the number of calls in groups ending because of the type (1) update and the number of calls in groups ending because of the type (2) update.
Because each type (1) update halves the value of $\epsilon_{k}$ and the algorithm terminates when $\epsilon_{k}\leq\epsilon,$ there can be at most $k_{1}=\lceil\log(\frac{2\epsilon_{0}}{\epsilon})\rceil=\lceil\log(\frac{2\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|}{\epsilon})\rceil$ type (1) updates. Because the updates of type (FixHal) satisfy $\|{\bm{y}}_{j}-{\bm{x}}_{*}\|\leq\|{\bm{y}}_{0}-{\bm{x}}_{*}\|\leq\|{\bm{x}}_{0}-{\bm{x}}_{*}\|,$ it follows by the triangle inequality that $\|{\bm{y}}_{j}-{\bm{y}}_{0}\|\leq 2\|{\bm{x}}_{0}-{\bm{x}}_{*}\|$ at any point in the algorithm execution. As a consequence, the estimate $D$ can never be increased beyond $4\|{\bm{x}}_{0}-{\bm{x}}_{*}\|.$ Let us index the type (1) updates by $i=\{1,2,\dots,k_{1}\}$. Leading to the $i^{\mathrm{th}}$ type (1) update, the value of $\lambda_{k}$ satisfies $\lambda_{k}=\Omega(\frac{\epsilon_{i}}{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|})=\Omega(\frac{\epsilon_{0}/2^{i}}{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|})$, as $D={\mathcal{O}}(\|{\bm{x}}_{0}-{\bm{x}}_{*}\|)$ and the value of $\epsilon_{k}$ only gets updated on a type (1) update. It follows by Lemma 1 that the corresponding number of oracle calls is ${\mathcal{O}}(\frac{1}{\epsilon_{0}/(2^{i}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|)+\ln(1/\gamma)})$. Thus, following the same argument as in the proof of Proposition 1, the total number of oracle calls in groups leading to the type (1) updates is
${\mathcal{O}}\Big{(}\min\Big{\{}\ln\Big{(}\frac{\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|}{\epsilon}\Big{)}\frac{1+\ln(1/\gamma)}{\ln(1/\gamma)},\,\ln(1/\gamma)+\frac{\ln(\gamma\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|/\epsilon)}{\ln(1/\gamma)},\,\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{\epsilon}\Big{\}}\Big{)}.$ (15)
It remains to bound the number of calls to ${\bm{T}}$ in the groups leading to the type (2) updates. A crucial observation here is that for each such group the number of oracle calls cannot be higher than ${\mathcal{O}}(\frac{1}{\lambda_{k}+\ln(1/\gamma)})$ for the current estimate of $\lambda_{k},$ because otherwise the fixed-point error would halve and we would reach a type (1) update. The total number of type (2) update groups is $k_{2}=\lceil\log(\frac{4\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{D_{0}})\rceil=\lceil\log(\frac{4\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|})\rceil,$ since each type (2) update leads to a doubling of the estimate $D,$ which is initialized at $D_{0}=\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|$ and bounded by $4\|{\bm{x}}_{0}-{\bm{x}}_{*}\|,$ as argued above. Let us index the groups leading to type (2) updates by $i=\{1,2,\dots,k_{2}\}.$ In the $i^{\mathrm{th}}$ group, the value of corresponding $\lambda_{k}$ is bounded below by $\Omega(\frac{\epsilon}{2^{i}D_{0}}),$ as $\epsilon_{k}$ cannot be reduced below $\epsilon/2$ and each type (2) update doubles the estimate $D.$ Following the argument from the proof of Proposition 1, we can thus bound the total number of oracle calls corresponding to type (2) updates by
${\mathcal{O}}\Big{(}\min\Big{\{}\ln\Big{(}\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|}\Big{)}\frac{1+\ln(1/\gamma)}{\ln(1/\gamma)},\,\ln(1/\gamma)+\frac{\ln(\gamma\|{\bm{x}}_{0}-{\bm{x}}_{*}\|/\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|)}{\ln(1/\gamma)},\,\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{\epsilon}\Big{\}}\Big{)}.$ (16)
The claimed bound on the oracle calls now follows by summing the bounds in Equation 15 and Equation 16. ∎
3.3 Comparison to Restarting-Based Approaches
We briefly discuss here how our results compare to restarting-based strategies as in prior work [Dia20] that sought to obtain an algorithm that simultaneously addresses cases where ${\bm{T}}$ can be nonexpansive or contractive, adapting to the better between the sublinear convergence rate (of Halpern iteration) and linear convergence rate (of Halpern iteration). While the results in [Dia20] apply only to Euclidean spaces, the underlying restarting idea can be applied more broadly to the classical Halpern iteration and any related algorithm that leads to results of the form [^8] Here we do not explicitly discuss the Krasnosel’skii-Mann iteration, as its (tight) convergence rate is $1/\sqrt{k}$ [BB96, CSV14, DW22], which is slower than the rate $1/k$ of Halpern iteration and leads to worse complexity results.
$\|{\bm{T}}({\bm{x}}_{k})-{\bm{x}}_{k}\|={\mathcal{O}}\Big{(}\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{k}\Big{)},$ (17)
after running the algorithm for $k\geq 1$ iterations, where the Lipschitz constant $\gamma$ of ${\bm{T}}$ satisfies $\gamma\leq 1$ and ${\bm{x}}_{*}$ is a fixed point of ${\bm{T}}$ (assuming it exists).
When $\gamma<1,$ since ${\bm{T}}({\bm{x}}_{*})={\bm{x}}_{*}$, we also have, by the triangle inequality, that $\|{\bm{x}}-{\bm{x}}_{*}\|\leq\|{\bm{x}}-{\bm{T}}({\bm{x}})\|+\|{\bm{T}}({\bm{x}})-{\bm{T}}({\bm{x}}_{*})\|\leq\|{\bm{x}}-{\bm{T}}({\bm{x}})\|+\gamma\|{\bm{x}}-{\bm{x}}_{*}\|$ and so $\|{\bm{x}}-{\bm{x}}_{*}\|\leq\frac{\|{\bm{x}}-{\bm{T}}({\bm{x}})\|}{1-\gamma}$. As a consequence, the bound in (17) implies $\|{\bm{T}}({\bm{x}}_{k})-{\bm{x}}_{k}\|={\mathcal{O}}\Big{(}\frac{\|{\bm{x}}_{0}-{\bm{T}}({\bm{x}}_{0})\|}{(1-\gamma)k}\Big{)}$, and so the fixed-point error $\|{\bm{T}}({\bm{x}}_{k})-{\bm{x}}_{k}\|$ can be halved within $k={\mathcal{O}}(1/(1-\gamma))$ iterations. Thus, by restarting the algorithm each time the fixed-point error halves, we get an algorithm that converges to a fixed-point error $\epsilon>0$ within ${\mathcal{O}}\big{(}\frac{\ln(\|{\bm{T}}({\bm{x}}_{0}-{\bm{x}}_{0}\|/\epsilon)}{1-\gamma}\big{)}$ iterations. When $\gamma$ is close to one, $\ln(1/\gamma)\approx 1-\gamma$ and this bound matches the optimal iteration complexity ${\mathcal{O}}\big{(}\frac{\ln(\|{\bm{T}}({\bm{x}}_{0}-{\bm{x}}_{0}\|/\epsilon)}{\ln(1/\gamma)}\big{)}$ of Picard iteration. However, when $\gamma$ is small, this bound is suboptimal, as $\ln(1/\gamma)$ can be much larger than $1-\gamma.$ Observe that the oracle complexity of GHAL and AdaGHAL in this case is ${\mathcal{O}}\big{(}\ln(1/\gamma)+\frac{\ln(\gamma\|{\bm{T}}({\bm{x}}_{0}-{\bm{x}}_{0}\|/\epsilon)}{\ln(1/\gamma)}\big{)}$, which is lower than ${\mathcal{O}}\big{(}\frac{\ln(\|{\bm{T}}({\bm{x}}_{0}-{\bm{x}}_{0}\|/\epsilon)}{1-\gamma}\big{)}$ whenever $\gamma$ is small and closer to the optimal iteration complexity of Picard iteration, matching it up to constant factors whenever $\gamma=\Omega(\sqrt{\epsilon/\left\lVert{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\right\rVert})$. This complexity advantage is clearly observed in the numerical examples in Section 5.
We remark here that (as also observed in [Dia20]) the oracle complexity bound of ${\mathcal{O}}\big{(}\frac{\ln(\|{\bm{T}}({\bm{x}}_{0}-{\bm{x}}_{0}\|/\epsilon)}{\mu}\big{)}$ for restart-based strategies applies even for operators that are not contractive, but instead are nonexpansive and satisfy a local error bound of the form $\|{\bm{T}}({\bm{x}})-{\bm{x}}\|\geq\mu\,\mathrm{dist}({\bm{x}},{\mathcal{X}}_{*})$, where ${\mathcal{X}}_{*}$ denotes the set of fixed points of ${\bm{T}}$ and $\mathrm{dist}({\bm{x}},{\mathcal{X}}_{*})=\inf_{{\bm{y}}\in{\mathcal{X}}_{*}}\|{\bm{x}}-{\bm{y}}\|$. In this case, it can be argued that restarted (FixHal) with an appropriate choice of the step size $\lambda$ converges linearly.
Lemma 4.
Suppose that ${\bm{T}}$ is a nonexpansive operator that satisfies $\|{\bm{T}}({\bm{x}})-{\bm{x}}\|\geq\mu\,\mathrm{dist}({\bm{x}},{\mathcal{X}}_{*})$, where ${\mathcal{X}}_{*}$ denotes the set of fixed points of ${\bm{T}}$ (assuming ${\mathcal{X}}_{*}$ is non-empty). Let ${\bm{x}}_{k}$ denote the $k^{\rm th}$ iterate of (FixHal) applied to ${\bm{T}},$ initialized at some ${\bm{x}}_{0}$ and using some step size $\lambda\in(0,1)$. Let $\beta\in(0,1).$ If $\lambda\leq\frac{\mu\beta/4}{1+\mu\beta/4},$ then $\|{\bm{T}}({\bm{x}}_{k})-{\bm{x}}_{k}\|\leq\beta\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|$ after at most $k=\lceil\frac{\ln(2/\beta)}{\ln(1/(1-\lambda))}\rceil$ iterations. As a result, choosing $\beta=1/2$ and restarting (FixHal) each time the fixed-point error is halved, we have that $\|{\bm{T}}({\bm{x}}_{k})-{\bm{x}}_{k}\|\leq\epsilon$ using a total of ${\mathcal{O}}\big{(}\frac{\ln(\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|/\epsilon)}{\mu}\big{)}$ oracle queries to ${\bm{T}}$ (equivalently, iterations of (FixHal)).
Proof.
From Lemma 1, we have that for all $k\geq 1$ and all ${\bm{x}}_{*}\in{\mathcal{X}}_{*},$ $\|{\bm{T}}({\bm{x}}_{k})-{\bm{x}}_{k}\|\leq(1-\lambda)^{k}\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|+\frac{2\lambda}{1-\lambda}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|$. Thus,
$\|{\bm{T}}({\bm{x}}_{k})-{\bm{x}}_{k}\|$ $\leq(1-\lambda)^{k}\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|+\frac{2\lambda}{1-\lambda}\mathrm{dist}({\bm{x}}_{0},{\mathcal{X}}_{*})$
$\leq(1-\lambda)^{k}\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|+\frac{2\lambda}{1-\lambda}\frac{1}{\mu}\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|,$
where the last inequality is by the lemma assumption. The first lemma claim now follows by solving the inequalities $(1-\lambda)^{k}\leq\beta/2$ and $\frac{2\lambda}{1-\lambda}\frac{1}{\mu}\|{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\|\leq\beta/2$ for $k$ and $\lambda,$ respectively.
Now consider restarting (FixHal) each time the fixed-point error of the iterate ${\bm{x}}_{k}$ gets halved (i.e., whenever $\left\lVert{\bm{T}}({\bm{x}}_{k})-{\bm{x}}_{k}\right\rVert\leq(1/2)\left\lVert{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\right\rVert$). Here, restarting means that the new calls to the iteration (FixHal) are initialized at the output ${\bm{x}}_{k}$ that triggered the restart. Using the first part of the lemma, a restart is triggered every $k=\lceil\frac{\ln(1/(2\beta))}{\ln(1/(1-\lambda))}\rceil={\mathcal{O}}\big{(}\frac{1}{\mu}\big{)}$ iterations, in which case the fixed-point error is halved. Hence, there can be at most $\log_{2}\lceil\left\lVert 2{\bm{T}}({\bm{x}}_{0})-{\bm{x}}_{0}\right\rVert/\epsilon\rceil$ restarts in total, proving the second claim. ∎
We remark here that the result of Lemma 4 does not directly translate into linear convergence for GHAL and AdaGHAL under a local error bound, as both algorithms progressively decrease the step size $\lambda$ as the fixed-point error of the iterates is decreased. While it is possible to change the while-loop conditions of these two algorithms to prevent reducing the step size once it is sufficiently small, this would, however, come at a cost of higher ${\mathcal{O}}\big{(}\frac{\ln(\|{\bm{T}}({\bm{x}}_{0}-{\bm{x}}_{0}\|/\epsilon)}{1-\gamma}\big{)}$ oracle complexity for contractive operators with a small value of $\gamma.$ For this reason, we do not consider such algorithm modifications and defer investigation of convergence under local error bounds to future work.
4 Gradually Expansive Operators and Convergence of GHAL
Even though improving the results under mild operator expansion seems unlikely, we argue below that by using Algorithm 1, we can ensure convergence to a fixed point for a class of operators with gradual expansion (defined below), whose bound on the Lipschitz constant is dependent on the fixed-point error (thus the bound is required to gradually decrease with the fixed-point error). The basic idea is that in Algorithm 1, the inner while loop can ensure a reduction in error within a finite number of iterations as long as $(1-\lambda_{k})\left\lVert{\bm{T}}({\bm{y}}_{j+1})-{\bm{T}}({\bm{y}}_{j})\right\rVert<\left\lVert{\bm{y}}_{j+1}-{\bm{y}}_{j}\right\rVert.$ The below definition of gradually expansive operators was obtained by ensuring such a condition could be satisfied.
Definition 1 ($\alpha$-Gradually Expansive Operators).
Given an operator ${\bm{T}}$ that satisfies 1, we say that ${\bm{T}}$ is $\alpha$-gradually expansive for some $\alpha\in[0,1)$, if the following holds: for all ${\bm{x}},{\bm{y}}\in{\mathcal{C}}$ and all $\epsilon>0$,
$\max\{\|{\bm{T}}({\bm{x}})-{\bm{x}}\|,\|{\bm{T}}({\bm{y}})-{\bm{y}}\|\}\leq\epsilon\;\Rightarrow\;\|{\bm{T}}({\bm{x}})-{\bm{T}}({\bm{y}})\|\leq\Big{(}1+\frac{\alpha\epsilon}{D}\Big{)}\|{\bm{x}}-{\bm{y}}\|.$
It is immediate that the above definition generalizes nonexpansive operators, as the case $\alpha=0$ corresponds to nonexpansivity. However, for $\alpha>0,$ the operator need not be nonexpansive, with allowed expansion depending on the fixed-point error relative to the diameter of the set ${\mathcal{C}}.$
Below, we provide convergence results by considering the setting in which $D$ is known, so Algorithm 1 applies. We note here that although it seems plausible that Algorithm 2 could also be applied to this setting (with a suitable safeguard as in Algorithm 1), we do not consider it for two reasons: (1) when the considered operator is potentially expansive, it becomes challenging to even guarantee that the fixed-point error exists (in finite-dimensional settings) or is finitely bounded (in inifinte-dimensional settings; recall the discussion from Section 2.2); and (2) handling unknown diameter of the set would further complicate an already fairly technical argument for limited returns, given (1).
Our main result is summarized in the following theorem.
Theorem 2.
Let ${\bm{T}}:{\mathcal{C}}\to{\mathcal{C}}$ be an $\alpha$-gradually expansive operator for $\alpha\in(0,\sqrt{2}-1)$, where the diameter of ${\mathcal{C}}$ is bounded by $D<\infty.$ Then, given $\epsilon>0,$ if Algorithm 1 is invoked with parameters $\beta\in(0,1)$ such that $\beta^{2}>\alpha\frac{1+\alpha}{1-\alpha}$ and $\beta^{\prime}\in[\frac{1}{1+c},1)$, then it outputs $\hat{\bm{x}}_{k}\in{\mathcal{C}}$ such that $\|{\bm{T}}(\hat{\bm{x}}_{k})-\hat{\bm{x}}_{k}\|\leq\epsilon$ using ${\mathcal{O}}\Big{(}\ln\Big{(}\frac{1}{\beta(1-\beta)}\Big{)}\min\Big{\{}\frac{D}{c\beta\epsilon},\frac{D^{2}}{\beta^{2}\epsilon^{2}}\Big{\}}\Big{)}$ oracle queries to ${\bm{T}},$ where $c>0$ is the largest constant such that $\beta^{2}=\frac{(1+c)\alpha(1+\alpha)}{1-(1+c)\alpha}$.
In particular, if $\alpha\in(0,0.4],$ then Algorithm 1 applied with $\beta=0.975$, $\beta^{\prime}=\frac{1}{1.01}$ outputs $\hat{\bm{x}}_{k}$ with $\|{\bm{T}}(\hat{\bm{x}}_{k})-\hat{\bm{x}}_{k}\|\leq\epsilon$ using ${\mathcal{O}}\big{(}\frac{D}{\epsilon}\big{)}$ oracle queries to ${\bm{T}}.$
Proof.
We first note that the parameter $\beta^{\prime}$ was chosen so that the safeguard condition in Algorithm 1 never gets triggered, as will be argued below by proving that all iterates ${\bm{y}}_{j}$ satisfy $\|{\bm{y}}_{j+1}-{\bm{y}}_{j}\|<(1-\beta^{\prime}\lambda_{k})\|{\bm{y}}_{j}-{\bm{y}}_{j-1}\|$ for $j\geq 2$ (see (23) in the proof below). Second, because each iteration of the while loop in Algorithm 1 reduces the value of $\epsilon_{k}$ by a factor $\beta,$ there can be at most $K=1+\big{\lceil}\log_{1/\beta}\big{(}\frac{\epsilon_{0}}{\epsilon}\big{)}\big{\rceil}$ iterations before the while loop halts, at which point it must be $\|{\bm{T}}(\hat{\bm{x}}_{K})-\hat{\bm{x}}_{K}\|\leq\epsilon$ by its exit condition.
To bound the total number of oracle queries to ${\bm{T}},$ we need to bound the number of calls to (FixHal) in the inner while loop of Algorithm 1. Consider the $k^{\mathrm{th}}$ iteration of the outer while loop for $k\geq 1.$ Our goal is to bound the maximum number of inner while loop iterations $j$ that ensure $\|{\bm{T}}({\bm{y}}_{j})-{\bm{y}}_{j}\|\leq\epsilon_{k}.$ We argue first by induction on $j$ that for all $j\geq 0,$ $\|{\bm{T}}({\bm{y}}_{j})-{\bm{y}}_{j}\|\leq(1-\lambda_{k})(1+\alpha+\beta^{2})\epsilon_{k-1}.$ We then use this bound to conclude that by Definition 1, $\|{\bm{T}}({\bm{y}}_{j+1})-{\bm{T}}({\bm{y}}_{j})\|\leq(1+\frac{\alpha(1-\lambda_{k})(1+\alpha+\beta^{2})\epsilon_{k-1}}{D})\|{\bm{y}}_{j+1}-{\bm{y}}_{j}\|.$
For the base case, we consider $j=\{0,1\}.$ Because ${\bm{y}}_{0}=\hat{\bm{x}}_{k-1},$ we have from Algorithm 1 that $\|{\bm{T}}({\bm{y}}_{0})-{\bm{y}}_{0}\|\leq\epsilon_{k-1}.$ By the definition of ${\bm{y}}_{1}:=\lambda_{k}{\bm{y}}_{0}+(1-\lambda_{k}){\bm{T}}({\bm{y}}_{0})$ we then also have $\|{\bm{y}}_{1}-{\bm{y}}_{0}\|=(1-\lambda_{k})\|{\bm{T}}({\bm{y}}_{0})-{\bm{y}}_{0}\|\leq(1-\lambda_{k})\epsilon_{k-1}.$ To finish the base case, it remains to bound $\|{\bm{T}}({\bm{y}}_{1})-{\bm{y}}_{1}\|.$ By the definition of ${\bm{y}}_{1},$ subtracting ${\bm{T}}({\bm{y}}_{1})$ and using the triangle inequality,
$\|{\bm{T}}({\bm{y}}_{1})-{\bm{y}}_{1}\|$ $\leq\lambda_{k}\|{\bm{y}}_{0}-{\bm{T}}({\bm{y}}_{1})\|+(1-\lambda_{k})\|{\bm{T}}({\bm{y}}_{1})-{\bm{T}}({\bm{y}}_{0})\|$ (18)
$\leq(1-\lambda_{k})\beta^{2}\epsilon_{k-1}+(1-\lambda_{k})\|{\bm{T}}({\bm{y}}_{1})-{\bm{T}}({\bm{y}}_{0})\|,$
where we have used the definition of $\lambda_{k},$ by which $\frac{\lambda_{k}}{1-\lambda_{k}}=\beta^{2}\epsilon_{k-1}/D$ and the fact that $\|{\bm{T}}({\bm{y}}_{1})-{\bm{y}}_{0}\|\leq D.$ Observe that for any ${\bm{y}}\in{\mathcal{C}},$ $\|{\bm{T}}({\bm{y}})-{\bm{y}}\|\leq D,$ and thus for any ${\bm{x}},{\bm{y}}\in{\mathcal{C}},$ by Definition 1, we have $\|{\bm{T}}({\bm{x}})-{\bm{T}}({\bm{y}})\|\leq(1+\alpha)\|{\bm{x}}-{\bm{y}}\|.$ As a consequence, applying Definition 1, we have
$\|{\bm{T}}({\bm{y}}_{1})-{\bm{y}}_{1}\|$ $\leq(1-\lambda_{k})\beta^{2}\epsilon_{k-1}+(1-\lambda_{k})(1+\alpha)\|{\bm{y}}_{1}-{\bm{y}}_{0}\|$
$\leq(1-\lambda_{k})\beta^{2}\epsilon_{k-1}+(1-\lambda_{k})^{2}(1+\alpha)\epsilon_{k-1}$
$<(1-\lambda_{k})(1+\alpha+\beta^{2})\epsilon_{k-1}.$ (19)
Now suppose that for some $j\geq 1$ we have for all $j^{\prime}\in\{0,1,\dots,j\},$ $\|{\bm{T}}({\bm{y}}_{j^{\prime}})-{\bm{y}}_{j^{\prime}}\|<(1-\lambda_{k})(1+\alpha+\beta^{2})\epsilon_{k-1}$. Then, by Definition 1, for all $j^{\prime}\in\{0,1,\dots,j\},$
$\|{\bm{T}}({\bm{y}}_{j^{\prime}})-{\bm{T}}({\bm{y}}_{j^{\prime}-1})\|<\Big{(}1+\frac{\alpha(1+\alpha+\beta^{2})\epsilon_{k-1}}{D}\Big{)}\|{\bm{y}}_{j^{\prime}}-{\bm{y}}_{j^{\prime}-1}\|,$
which further implies that
$\|{\bm{y}}_{j^{\prime}+1}-{\bm{y}}_{j^{\prime}}\|$ $=(1-\lambda_{k})\|{\bm{T}}({\bm{y}}_{j^{\prime}})-{\bm{T}}({\bm{y}}_{j^{\prime}-1})\|$
$<(1-\lambda_{k})\Big{(}1+\frac{\alpha(1+\alpha+\beta^{2})\epsilon_{k-1}}{D}\Big{)}\|{\bm{y}}_{j^{\prime}}-{\bm{y}}_{j^{\prime}-1}\|.$
Observing that
$(1-\lambda_{k})\Big{(}1+\frac{\alpha(1+\alpha+\beta^{2})\epsilon_{k-1}}{D}\Big{)}$ $=\frac{1}{1+\beta^{2}\epsilon_{k-1}/D}\Big{(}1+\frac{\alpha(1+\alpha+\beta^{2})\epsilon_{k-1}}{D}\Big{)}\leq 1,$
where the inequality is from the condition $\beta^{2}\geq\alpha\frac{1+\alpha}{1-\alpha}$ from the theorem statement, we conclude that for all $j^{\prime}\in\{0,1,\dots,j\},$ we have $\|{\bm{y}}_{j^{\prime}+1}-{\bm{y}}_{j^{\prime}}\|<\|{\bm{y}}_{j^{\prime}}-{\bm{y}}_{j^{\prime}-1}\|$ and so it must be
$\|{\bm{y}}_{j+1}-{\bm{y}}_{j}\|<\|{\bm{y}}_{1}-{\bm{y}}_{0}\|\leq(1-\lambda_{k})\epsilon_{k-1}.$ (20)
Using the definition of ${\bm{y}}_{j+1}$ and the triangle inequality,
$\|{\bm{T}}({\bm{y}}_{j+1})-{\bm{y}}_{j+1}\|$ $\leq\lambda_{k}\|{\bm{y}}_{0}-{\bm{T}}({\bm{y}}_{j+1})\|+(1-\lambda_{k})\|{\bm{T}}({\bm{y}}_{j+1})-{\bm{T}}({\bm{y}}_{j})\|$
$<(1-\lambda_{k})\beta^{2}\epsilon_{k-1}+(1+\alpha)(1-\lambda_{k})^{2}\epsilon_{k-1}$
$<(1-\lambda_{k})(1+\alpha+\beta^{2})\epsilon_{k-1},$
where the second inequality uses the definition of $\lambda_{k}$, $\|{\bm{y}}_{0}-{\bm{T}}({\bm{y}}_{j+1})\|\leq D,$ $\|{\bm{T}}({\bm{y}}_{j+1})-{\bm{T}}({\bm{y}}_{j})\|\leq(1+\alpha)\|{\bm{y}}_{j+1}-{\bm{y}}_{j}\|$ and Equation 20, while the third inequality uses $\lambda_{k}\in(0,1).$
Thus, we conclude by induction on $j$ that all iterates $j$ in the inner while loop of Algorithm 1 satisfy $\|{\bm{T}}({\bm{y}}_{j})-{\bm{y}}_{j}\|<(1-\lambda_{k})(1+\alpha+\beta^{2})\epsilon_{k-1}.$ As a consequence, for all $j\geq 0,$
$\|{\bm{T}}({\bm{y}}_{j+1})-{\bm{T}}({\bm{y}}_{j})\|\leq\Big{(}1+(1-\lambda_{k})\frac{\alpha\epsilon_{k-1}}{D}(1+\alpha+\beta^{2})\Big{)}\|{\bm{y}}_{j+1}-{\bm{y}}_{j}\|.$ (21)
Suppose now that $\beta^{2}=(1+c)\alpha(1+\alpha+\beta^{2})$ for some $c>0$ (in other words, $\beta^{2}=\frac{(1+c)\alpha(1+\alpha)}{1-(1+c)\alpha}$), where we require $\beta\in(0,1)$ (note that such a choice must exist, due to the assumption that $\alpha<\sqrt{2}-1$). Then:
$1-\lambda_{k}$ $=\frac{1}{1+\beta^{2}\epsilon_{k-1}/D}$
$=\frac{1}{1+(1+c)\alpha(1+\alpha+\beta^{2})\epsilon_{k-1}/D}.$
A rearrangement of the last equality gives
$\alpha(1+\alpha+\beta^{2})\epsilon_{k-1}/D=\frac{1}{1+c}\cdot\frac{\lambda_{k}}{1-\lambda_{k}},$
which, when plugged back into (21) leads to
$\|{\bm{T}}({\bm{y}}_{j+1})-{\bm{T}}({\bm{y}}_{j})\|\leq\Big{(}1+\frac{\lambda_{k}}{1+c}\Big{)}\|{\bm{y}}_{j+1}-{\bm{y}}_{j}\|,\;\forall j\geq 0.$ (22)
As a consequence, we have that for all $j\geq 1,$
$\|{\bm{y}}_{j+1}-{\bm{y}}_{j}\|\leq\Big{(}1-\frac{\lambda_{k}}{1+c}(c+\lambda_{k})\Big{)}\|{\bm{y}}_{j}-{\bm{y}}_{j-1}\|$ (23)
Applying Inequality (23) recursively and recalling that $\|{\bm{y}}_{1}-{\bm{y}}_{0}\|=(1-\lambda_{k})\|{\bm{T}}({\bm{y}}_{0})-{\bm{y}}_{0}\|\leq(1-\lambda_{k})\epsilon_{k-1},$ we get that for all $j\geq 1,$
$\|{\bm{y}}_{j+1}-{\bm{y}}_{j}\|\leq(1-\lambda_{k})\Big{(}1-\frac{\lambda_{k}}{1+c}(c+\lambda_{k})\Big{)}^{j}\epsilon_{k-1}$ (24)
Combining Inequality (24) with $\|{\bm{T}}({\bm{y}}_{j})-{\bm{y}}_{j}\|\leq\frac{1}{1-\lambda_{k}}\|{\bm{y}}_{j+1}-{\bm{y}}_{j}\|+\frac{\lambda_{k}}{1-\lambda_{k}}\|{\bm{y}}_{0}-{\bm{y}}_{j}\|$ (see the proof of Lemma 1), we conclude that for all $j\geq 1,$
$\|{\bm{T}}({\bm{y}}_{j})-{\bm{y}}_{j}\|$ $\leq\Big{(}1-\frac{\lambda_{k}}{1+c}(c+\lambda_{k})\Big{)}^{j}\epsilon_{k-1}+\frac{\lambda_{k}}{1-\lambda_{k}}\|{\bm{y}}_{0}-{\bm{y}}_{j}\|$
$\leq\Big{(}1-\frac{\lambda_{k}}{1+c}(c+\lambda_{k})\Big{)}^{j}\epsilon_{k-1}+\beta\epsilon_{k}.$ (25)
Because $\big{(}1-\frac{\lambda_{k}}{1+c}(c+\lambda_{k})\big{)}^{j}\epsilon_{k-1}\leq(1-\beta)\epsilon_{k}$ for all $j\geq\frac{2(1+c)\ln(1/(\beta(1-\beta)))}{\lambda_{k}(c+\lambda_{k})}$, we conclude from Inequality (25) that the maximum number of oracle queries to ${\bm{T}}$ that the inner while loop of Algorithm 1 makes in iteration $k$ of the outer while loop is $j_{k}=\big{\lceil}\frac{2(1+c)\ln(1/(\beta(1-\beta)))}{\lambda_{k}(c+\lambda_{k})}\big{\rceil}$.
To complete bounding the total number of oracle queries to ${\bm{T}},$ it remains to sum $j_{k}$ over $k\in\{1,2,\dots,K\}$. Doing so leads to the bound on the total number of queries given by
$\sum_{k=1}^{K}j_{k}$ $={\mathcal{O}}\Big{(}\ln\Big{(}\frac{1}{\beta(1-\beta)}\Big{)}\sum_{k=1}^{1+\lceil\log_{1/\beta}(\frac{\epsilon_{0}}{\epsilon})\rceil}\frac{1}{\lambda_{k}(c+\lambda_{k})}\Big{)}$
$={\mathcal{O}}\Big{(}\ln\Big{(}\frac{1}{\beta(1-\beta)}\Big{)}\sum_{k=1}^{1+\lceil\log_{1/\beta}(\frac{\epsilon_{0}}{\epsilon})\rceil}\min\Big{\{}\frac{D}{c\beta\epsilon_{k}},\frac{D^{2}}{\beta^{2}\epsilon_{k}^{2}}\Big{\}}\Big{)}$
$={\mathcal{O}}\Big{(}\ln\Big{(}\frac{1}{\beta(1-\beta)}\Big{)}\sum_{k=1}^{1+\lceil\log_{1/\beta}(\frac{\epsilon_{0}}{\epsilon})\rceil}\min\Big{\{}\frac{D}{c\beta\epsilon_{0}\beta^{k}},\frac{D^{2}}{\beta^{2}\epsilon_{0}^{2}\beta^{2k}}\Big{\}}\Big{)}$
$={\mathcal{O}}\Big{(}\ln\Big{(}\frac{1}{\beta(1-\beta)}\Big{)}\min\Big{\{}\frac{D}{c\beta\epsilon},\frac{D^{2}}{\beta^{2}\epsilon^{2}}\Big{\}}\Big{)},$
as claimed. For the second part of the claim, suppose that $\alpha\in(0,0.4].$ Then it is not hard to check (e.g., using numerical solvers or solving by hand) that $\beta=0.975$ satisfies $\beta^{2}=(1+c)\alpha(1+\alpha+\beta^{2})$ for $c\geq 0.01.$ As a consequence, since $c,\beta$ are both absolute constants in this case, we get that the upper bound on the number of oracle queries to ${\bm{T}}$ is ${\mathcal{O}}\big{(}\frac{D}{\epsilon}\big{)}.$
An interesting consequence of Theorem 2 is that it implies that for $\alpha$-gradually expansive operators with $\alpha\in[0,\sqrt{2}-1),$ we can guarantee that $\inf_{{\bm{x}}\in{\mathcal{C}}}\left\lVert{\bm{T}}({\bm{x}})-{\bm{x}}\right\rVert=0,$ even in infinite-dimensional spaces, despite the Lipschitz constant of ${\bm{T}}$ being possibly as large as $\gamma=1+\alpha\in[1,\sqrt{2})$. Recall here that, without additional assumptions, in infinite-dimensional Banach spaces, the best that can be shown for a $\gamma$-Lipschitz operator ${\bm{T}}$ is that $\inf_{{\bm{x}}\in{\mathcal{C}}}\left\lVert{\bm{T}}({\bm{x}})-{\bm{x}}\right\rVert\leq 1-1/\gamma,$ and this inequality is tight in general [Goe73].
4.1 A Simple Example and Properties
Gradual expansion is a technical condition that, as mentioned earlier, was discovered when studying the convergence properties of GHAL. However, its interpretation is intuitive: it bounds the excess expansion ($\gamma-1$) of an operator between two points proportional to the maximum of fixed-point errors of the two points relative to the diameter of the set. One may wonder then what kind of operators could satisfy such a condition. In this subsection, we provide one such simple example as a possible starting point, while we leave their further study to future work.
Lemma 5.
Given $d\geq 1,$ $\alpha\in[0,1),$ $D>0$, and ${\mathcal{C}}=[-D/2,D/2]^{d},$ let ${\bm{T}}({\bm{x}}):=\mathrm{Proj}_{{\mathcal{C}}}({\bm{x}}+\exp(\alpha{\bm{x}}/D))$, where $\mathrm{Proj}_{{\mathcal{C}}}$ denotes orthogonal projection onto ${\mathcal{C}}$ and $\exp(\alpha{\bm{x}}/D)$ is applied coordinate-wise. Then ${\bm{T}}({\bm{x}})$ is $\alpha$-gradually expansive w.r.t. the $\ell_{\infty}$ norm.
Proof.
Observe that ${\bm{T}}:{\mathcal{C}}\to{\mathcal{C}}$, by its definition. The claim of the lemma is immediate when $\alpha=0$, as ${\bm{T}}$ is nonexpansive in this case (coordinate-wise shift + clipping to the interval $[-D/2,D/2]$), so assume that $\alpha\in(0,1).$ Consider first a univariate operator $t(x)$ defined via $\mathrm{Proj}_{[-D/2,D/2]}(x+f(x))$, where $\mathrm{Proj}_{[-D/2,D/2]}(t)=\min\{D/2,\max\{-D/2,t\}\}$ is the projection onto the unit interval $[-D/2,D/2]$ and we define $f(x)=\exp(\alpha x/D)$. In the rest of the proof, we argue that $t$ is $\alpha$-gradually expansive. This suffices for proving that ${\bm{T}}$ is $\alpha$-gradually expansive, since (using notation (i) for indexing the coordinates of a vector)
$\|{\bm{T}}({\bm{x}})-{\bm{T}}({\bm{y}})\|_{\infty}$ $=\max_{1\leq i\leq d}|t({\bm{x}}^{(i)})-t({\bm{y}}^{(i)})|$
$\leq\max_{1\leq i\leq d}\Big{(}1+\frac{\alpha\max\{|t({\bm{x}}^{(i)})-{\bm{x}}^{(i)}|,\,|t({\bm{y}}^{(i)})-{\bm{y}}^{(i)}|\}}{D}\Big{)}|{\bm{x}}^{(i)}-{\bm{y}}^{(i)}|$
$\leq\Big{(}1+\frac{\alpha\max\{\left\lVert{\bm{T}}({\bm{x}})-{\bm{x}}\right\rVert_{\infty},\,\left\lVert{\bm{T}}({\bm{y}})-{\bm{y}}\right\rVert_{\infty}\}}{D}\Big{)}\left\lVert{\bm{x}}-{\bm{y}}\right\rVert_{\infty},$
where the first inequality is by $t$ being $\alpha$-gradually expansive.
Observe that $f^{\prime}(x)=\frac{\alpha}{D}f(x)$ and so $|f(x)-f(y)|\leq\max\{|f^{\prime}(x)|,|f^{\prime}(y)|\}|x-y|=\frac{\alpha}{D}\max\{|f(x)|,|f(y)|\}|x-y|$. Thus, it is immediate that $|t(x)-t(y)|\leq(1+\frac{\alpha}{D}\max\{|f(x)|,|f(y)|\})|x-y|$. As a consequence, if $t(x)=x+f(x)$ and $t(y)=y+f(y)$ (i.e., if neither $x$ nor $y$ gets clipped by the projection onto $[-D/2,D/2]$), then the inequality defining $\alpha$-gradually expansive operators applies.
Observe further that $t(x)$ is increasing in $x,$ so the only way it can happen that $t(x)\neq x+f(x)$ is if $x+f(x)>D/2,$ in which case $t(x)=D/2.$ If both $x$ and $y$ get clipped, then both are equal to $D/2,$ so in this case $|t(x)-t(y)|=0$ and the inequality defining $\alpha$-gradually expansive operators holds trivially. Thus, it remains to argue that this inequality still holds when only one of $x,y$ gets clipped. Suppose w.l.o.g. that $t(x)=D/2$ and $t(y)<D/2.$ Observe that in this case it must be $x>y$. Let $\Delta(x)=|t(x)-x|=t(x)-x=D/2-x.$ We now have:
$|t(x)-t(y)|$ $=D/2-y-e^{\alpha y/D}$
$=x-y+\Delta(x)-e^{\alpha y/D}.$
To complete the proof of the lemma, it suffices to argue that in this case it holds that
$\Delta(x)-e^{\alpha y/D}\leq\frac{\alpha}{D}\max\{\Delta(x),\,e^{\alpha y/D}\}(x-y),$ (26)
as $x>y,$ $\Delta(x)=|t(x)-x|$ and $e^{\alpha y/D}=|t(y)-y|$.
When $\Delta(x)\leq e^{\alpha y/D}$, (26) holds trivially, as the left-hand side is non-positive, while the right-hand side is positive. Thus, suppose now that $\Delta(x)>e^{\alpha y/D}$. Then, (26) is equivalent to
$\Delta(x)-e^{\alpha y/D}\leq\frac{\alpha}{D}\Delta(x)(x-y),$
which, rearranging and recalling that $\Delta(x)=D/2-x\geq 0,$ is equivalent to
$(D/2-x)\Big{(}1-\frac{\alpha}{D}(x-y)\Big{)}\leq e^{\alpha y/D}.$ (27)
Observe that the left-hand side of (27) is decreasing in $x.$ This is true because both $(D/2-x)$ and $\big{(}1-\frac{\alpha}{D}(x-y)\big{)}$ are non-negative and decreasing in $x$ (the latter term is non-negative because $(x-y)\leq D$ and $\alpha<1$). Thus, the left-hand side is maximized when $x$ is minimized, which for the present case means that $x$ is such that $x+e^{cx/D}=D/2$ (the smallest value at which $x$ gets clipped; if no such point exists, then $t(x)=D/2$ for all $x\in[-D/2,D/2]$ and the lemma claim holds trivially). Plugging this value of $x$ into (27), we get that it suffices to ensure that
$e^{cx/D}\Big{(}1-\frac{\alpha}{D}(x-y)\Big{)}\leq e^{\alpha y/D}.$
This inequality is equivalent to $1+\frac{\alpha}{D}(y-x)\leq e^{\alpha/D(y-x)}$, which is true, as $e^{a}\geq 1+a$ for any real number $a$. ∎
Lemma 5 provides an example of an operator ${\bm{T}}$ that is $\alpha$-gradually expansive (but not nonexpansive, for any $\alpha\in(0,1)$ and $D>2$). However, this example is not interesting for numerical evaluations, as any reasonable fixed-point iteration (including Picard) would converge to a fixed point of ${\bm{T}}$. This is because ${\bm{T}}({\bm{x}})\geq{\bm{x}}$ coordinate-wise, and as soon as any of the coordinates reaches $D/2,$ a fixed point of ${\bm{T}}$ is found. Nevertheless, we hope that this simple example can serve as a starting point for constructing more interesting gradually expansive operators, possibly by trying to replace the separable exponential mapping in its definition by a different mapping that would prevent Picard iteration from converging, while retaining the gradual expansion property. Towards this goal, we also state and prove a simple structural lemma that provides sufficient conditions for an operator to be $\alpha$-gradually expansive.
Lemma 6.
Let ${\bm{T}}:{\mathcal{C}}\to{\mathcal{C}}$ be an operator that satisfies 1. Let ${\bm{F}}:={\bm{T}}-{\bm{\mathrm{Id}}}.$ If for all ${\bm{x}},{\bm{y}}\in{\mathcal{C}}$ it holds that $\|{\bm{F}}({\bm{x}})-{\bm{F}}({\bm{y}})\|\leq\frac{\alpha}{D}\max\{\|{\bm{F}}({\bm{x}})\|,\,\|{\bm{F}}({\bm{y}})\|\}\|{\bm{x}}-{\bm{y}}\|$ for some $\alpha\in[0,1)$, then ${\bm{T}}$ is $\alpha$-gradually expansive. Further, in this case, the operator $\widetilde{\bm{T}}={\bm{\mathrm{Id}}}+{\bm{F}}\circ{\bm{R}}$ is $\alpha$-gradually expansive for any nonexpansive operator ${\bm{R}}:{\mathcal{C}}\to{\mathcal{C}}.$
Proof.
The first claim of the lemma is immediate, as, by the triangle inequality and the definition of ${\bm{F}},$ we have
$\|{\bm{T}}({\bm{x}})-{\bm{T}}({\bm{y}})\|$ $\leq\|{\bm{x}}-{\bm{y}}\|+\|{\bm{F}}({\bm{x}})-{\bm{F}}({\bm{y}})\|$
$\leq\Big{(}1+\frac{\alpha}{D}\max\{\|{\bm{F}}({\bm{x}})\|,\,\|{\bm{F}}({\bm{y}})\|\}\Big{)}\|{\bm{x}}-{\bm{y}}\|$
$=\Big{(}1+\frac{\alpha}{D}\max\{\|{\bm{T}}({\bm{x}})-{\bm{x}}\|,\,\|{\bm{T}}({\bm{y}})-{\bm{y}}\|\}\Big{)}\|{\bm{x}}-{\bm{y}}\|.$
For the remaining lemma claim, since $\|{\bm{F}}({\bm{x}})-{\bm{F}}({\bm{y}})\|\leq\frac{\alpha}{D}\max\{\|{\bm{F}}({\bm{x}})\|,\,\|{\bm{F}}({\bm{y}})\|\}\|{\bm{x}}-{\bm{y}}\|$ and ${\bm{R}}:{\mathcal{C}}\to{\mathcal{C}}$ is nonexpansive, we have
$\|{\bm{F}}({\bm{R}}({\bm{x}}))-{\bm{F}}({\bm{R}}({\bm{y}}))\|$ $\leq\frac{\alpha}{D}\max\{\|{\bm{F}}({\bm{R}}({\bm{x}}))\|,\,\|{\bm{F}}({\bm{R}}({\bm{y}}))\|\}\|{\bm{R}}({\bm{x}})-{\bm{R}}({\bm{y}})\|$
$\leq\frac{\alpha}{D}\max\{\|{\bm{F}}({\bm{R}}({\bm{x}}))\|,\,\|{\bm{F}}({\bm{R}}({\bm{y}}))\|\}\|{\bm{x}}-{\bm{y}}\|$
$=\frac{\alpha}{D}\max\{\|\widetilde{\bm{T}}({\bm{x}})-{\bm{x}}\|,\,\|\widetilde{\bm{T}}({\bm{y}})-{\bm{y}}\|\}\|{\bm{x}}-{\bm{y}}\|,$
where the second inequality is by the nonexpansiveness of ${\bm{R}}$ and the last inequality is by the definition of $\widetilde{\bm{T}}.$ To complete the proof, it remains to use the triangle inequality by which $\|\widetilde{\bm{T}}({\bm{x}})-\widetilde{\bm{T}}({\bm{y}})\|\leq\|{\bm{x}}-{\bm{y}}\|+\|{\bm{F}}({\bm{R}}({\bm{x}}))-{\bm{F}}({\bm{R}}({\bm{y}}))\|\leq\big{(}1+\frac{\alpha}{D}\max\{\|\widetilde{\bm{T}}({\bm{x}})-{\bm{x}}\|,\,\|\widetilde{\bm{T}}({\bm{y}})-{\bm{y}}\|\}\big{)}\|{\bm{x}}-{\bm{y}}\|$. ∎
4.2 Implications of Mild and Gradual Expansion on Root Finding Problems
[…]
${\bm{G}}_{{\bm{u}}}+{\bm{u}}={\bm{x}}.$ (28)
As a consequence, given any $\epsilon\geq 0,$ we have that $\|{\bm{G}}_{\bm{u}}\|\leq\epsilon$ if and only if $\|{\bm{J}}({\bm{x}})-{\bm{x}}\|\leq\epsilon.$ Hence we can reduce the problem of finding an $\epsilon$-approximate root of ${\bm{F}}$ (namely, a point ${\bm{x}}$ such that $\exists{\bm{G}}\in{\bm{F}}({\bm{x}})$ such that $\|{\bm{G}}\|\leq\epsilon$) to the problem of finding an $\epsilon$-approximate fixed point of ${\bm{J}}.$ Thus, in the following, we provide sufficient conditions for ${\bm{F}}$ that ensure ${\bm{J}}$ is either mildly or gradually expansive, so the associated root finding problems can be approximately solved by finding an approximate fixed point of ${\bm{J}}$.
First, we establish equivalence between $\gamma$-Lipschitzness of ${\bm{J}}$ and a condition on ${\bm{F}}$ stated in the following lemma, which in turn leads to a sufficient condition on ${\bm{F}}$ that leads to a mildly nonexpansive resolvent. This lemma is simple to prove and is provided for completeness.
Lemma 7.
Given ${\bm{F}}:{\mathcal{E}}\rightrightarrows{\mathcal{E}}$ such that $\mathrm{dom}({\bm{F}})\neq\emptyset,$ let ${\bm{J}}=({\bm{F}}+{\bm{\mathrm{Id}}})^{-1}.$ If for some $\mu\in[0,1)$, any ${\bm{u}},{\bm{v}}\in\mathrm{dom}({\bm{F}})$, and any ${\bm{G}}_{u}\in{\bm{F}}({\bm{u}}),{\bm{G}}_{\bm{v}}\in{\bm{F}}({\bm{v}}),$ we have
$\langle{\bm{G}}_{\bm{u}}-{\bm{G}}_{\bm{v}},{\bm{u}}-{\bm{v}}\rangle\geq-\mu\|{\bm{u}}-{\bm{v}}\|^{2},$ (29)
then ${\bm{J}}$ is $\frac{1}{1-\mu}$-Lipschitz. In particular, suppose that $\sup_{{\bm{x}},{\bm{y}}\in\mathrm{dom}({\bm{F}})}\|{\bm{x}}-{\bm{y}}\|=D.$ Then a sufficient condition for ${\bm{J}}$ to be $\gamma$-Lipschitz with $\gamma=1+\frac{\epsilon\beta}{D}$ for some $\epsilon>0,$ $\beta\in(0,1)$ is that the above inequality applies with $\mu=\frac{\beta\epsilon}{D+\beta\epsilon}.$
Proof.
The second part of the lemma is an immediate corollary of the first part, so we only prove the first statement. Fix any ${\bm{x}},{\bm{y}}\in{\mathcal{E}}$ and let ${\bm{u}}={\bm{J}}({\bm{x}}),$ ${\bm{v}}={\bm{J}}({\bm{y}})$. Observe that ${\bm{u}},{\bm{v}}\in\mathrm{dom}({\bm{F}})$. Using the definition of the resolvent, we further have ${\bm{x}}={\bm{G}}_{\bm{u}}+{\bm{u}}$ for some ${\bm{G}}_{\bm{u}}\in{\bm{F}}({\bm{u}})$ and ${\bm{y}}={\bm{G}}_{\bm{v}}+{\bm{v}}$ for some ${\bm{G}}_{\bm{v}}\in{\bm{F}}({\bm{v}}).$ Thus, we have that
$\|{\bm{u}}-{\bm{v}}\|^{2}=\langle{\bm{x}}-{\bm{y}},{\bm{u}}-{\bm{v}}\rangle-\langle{\bm{G}}_{\bm{u}}-{\bm{G}}_{\bm{v}},{\bm{u}}-{\bm{v}}\rangle.$ (30)
Since, by assumption, $\langle{\bm{G}}_{\bm{u}}-{\bm{G}}_{\bm{v}},{\bm{u}}-{\bm{v}}\rangle\geq-\mu\|{\bm{u}}-{\bm{v}}\|^{2},$ rearranging (30) and applying Cauchy-Schwarz inequality, we get
$(1-\mu)\|{\bm{u}}-{\bm{v}}\|^{2}\leq\langle{\bm{x}}-{\bm{y}},{\bm{u}}-{\bm{v}}\rangle\leq\|{\bm{x}}-{\bm{y}}\|\|{\bm{u}}-{\bm{v}}\|.$
Recalling that ${\bm{u}}={\bm{J}}({\bm{x}}),{\bm{v}}={\bm{J}}({\bm{y}}),$ this in turn implies that ${\bm{J}}$ is $\frac{1}{1-\mu}$-Lipschitz. ∎
One way of satisfying the conditions of Lemma 7 is when ${\bm{F}}$ is the sum of a single-valued operator ${\bm{G}}$ that satisfies (29) and a normal cone of a convex set of bounded diameter $D.$ For instance, such a condition would hold for weakly convex-weakly concave min-max optimization problems with a sufficiently small (order $\epsilon/D$) weak convexity parameter $\mu$, without imposing any additional assumptions such as the existence of weak solutions to the associated variational inequality problem [RLLY22]. [^9] Assuming that there exists a weak (or Minty) solution to the associated variational inequality is very restricting. For instance, for Lipschitz operators, such an assumption suffices for solving the associated problem with basic extragradient-type methods without any further assumptions on weak monotonicity as in (29).
Finally, it is possible to derive a sufficient condition for the resolvent of a possibly nonmonotone operator ${\bm{F}}$ to be $\alpha$-gradually expansive, as summarized in the following lemma.
Lemma 8.
Given ${\bm{F}}:{\mathcal{E}}\rightrightarrows{\mathcal{E}}$ such that $\mathrm{dom}({\bm{F}})\neq\emptyset$ and $\sup_{{\bm{x}},{\bm{y}}\in\mathrm{dom}({\bm{F}})}\|{\bm{x}}-{\bm{y}}\|=D,$ let ${\bm{J}}=({\bm{F}}+{\bm{\mathrm{Id}}})^{-1}.$ If for some $\mu\in[0,1)$, any ${\bm{u}},{\bm{v}}\in\mathrm{dom}({\bm{F}})$, and any ${\bm{G}}_{u}\in{\bm{F}}({\bm{u}}),{\bm{G}}_{\bm{v}}\in{\bm{F}}({\bm{v}}),$ we have
$\langle{\bm{G}}_{\bm{u}}-{\bm{G}}_{\bm{v}},{\bm{u}}-{\bm{v}}\rangle\geq-\frac{\tau_{{\bm{u}},{\bm{v}}}}{1+\tau_{{\bm{u}},{\bm{v}}}}\|{\bm{u}}-{\bm{v}}\|^{2},$ (31)
where $\tau_{{\bm{u}},{\bm{v}}}:=\frac{\alpha}{D}\max\{\|{\bm{G}}_{\bm{u}}\|,\|{\bm{G}}_{\bm{v}}\|\}$, then ${\bm{J}}$ is $\alpha$-gradually expansive.
Proof.
Similar to the proof of Lemma 7, fix any ${\bm{x}},{\bm{y}}\in{\mathcal{E}}$ and let ${\bm{u}}={\bm{J}}({\bm{x}}),$ ${\bm{v}}={\bm{J}}({\bm{y}})$. Recall that ${\bm{x}}={\bm{G}}_{\bm{u}}+{\bm{u}}$ for some ${\bm{G}}_{\bm{u}}\in{\bm{F}}({\bm{u}})$ and ${\bm{y}}={\bm{G}}_{\bm{v}}+{\bm{v}}$ for some ${\bm{G}}_{\bm{v}}\in{\bm{F}}({\bm{v}}).$ Then:
$\|{\bm{u}}-{\bm{v}}\|^{2}=\langle{\bm{x}}-{\bm{y}},{\bm{u}}-{\bm{v}}\rangle-\langle{\bm{G}}_{\bm{u}}-{\bm{G}}_{\bm{v}},{\bm{u}}-{\bm{v}}\rangle.$ (32)
Since, by assumption, $\langle{\bm{G}}_{\bm{u}}-{\bm{G}}_{\bm{v}},{\bm{u}}-{\bm{v}}\rangle\geq-\frac{\tau_{{\bm{u}},{\bm{v}}}}{1+\tau_{{\bm{u}},{\bm{v}}}}\|{\bm{u}}-{\bm{v}}\|^{2},$ rearranging (32) leads to
$\|{\bm{u}}-{\bm{v}}\|^{2}\leq(1+\tau_{{\bm{u}},{\bm{v}}})\langle{\bm{x}}-{\bm{y}},{\bm{u}}-{\bm{v}}\rangle.$
Thus, using the Cauchy-Schwarz inequality and simplifying,
$\|{\bm{u}}-{\bm{v}}\|\leq(1+\tau_{{\bm{u}},{\bm{v}}})\|{\bm{x}}-{\bm{y}}\|.$
To complete the proof, it remains to recall that ${\bm{u}}={\bm{J}}({\bm{x}}),$ ${\bm{v}}={\bm{J}}({\bm{y}}),$ ${\bm{G}}_{\bm{u}}={\bm{x}}-{\bm{J}}({\bm{x}}),$ and ${\bm{G}}_{\bm{v}}={\bm{y}}-{\bm{J}}({\bm{y}}).$ Plugging these identities into the last inequality leads to the conclusion that ${\bm{J}}$ is $\alpha$-gradually expansive, by definition. ∎
A few remarks are in order here. First, observe that it is not necessary for (31) to hold for all elements of ${\bm{F}}({\bm{u}})$ but only those selected by the resolvent; i.e., those that satisfy ${\bm{G}}_{\bm{x}}\in{\bm{F}}({\bm{J}}({\bm{x}}))$ and ${\bm{x}}={\bm{G}}_{\bm{x}}+{\bm{J}}({\bm{x}})$ (though it is not immediately clear how to enforce such a condition). Second, perhaps the most basic example of an operator ${\bm{F}}$ to consider is when it is expressible as the sum of a 1-Lipschitz-continuous [^10] Here, 1-Lipschitz is without loss of generality, as we can rescale ${\bm{F}}_{1}$ by its Lipschitz constant. operator ${\bm{F}}_{1}$ and the normal cone of a closed convex set ${\bm{C}}$ with a nonempty interior. Then, in the interior of the set ${\mathcal{C}},$ we have ${\bm{F}}={\bm{F}}_{1}$, which is single-valued and 1-Lipschitz-continuous. The condition (31) for ${\bm{x}},{\bm{y}}$ in the interior of ${\mathcal{C}}$ then becomes
$\langle{\bm{F}}({\bm{x}})-{\bm{F}}({\bm{y}}),{\bm{x}}-{\bm{y}}\rangle\geq-\frac{(\alpha/D)\max\{\left\lVert{\bm{F}}({\bm{x}})\right\rVert,\left\lVert{\bm{F}}({\bm{y}})\right\rVert\}}{1+(\alpha/D)\max\{\left\lVert{\bm{F}}({\bm{x}})\right\rVert,\left\lVert{\bm{F}}({\bm{y}})\right\rVert\}}\|{\bm{x}}-{\bm{y}}\|.$
Since it is possible, in general, for $\max\{\left\lVert{\bm{F}}({\bm{x}})\right\rVert,\left\lVert{\bm{F}}({\bm{y}})\right\rVert\}$ to take values of order-$D$ (or even larger, if ${\mathcal{C}}$ does not contain a root of ${\bm{F}}_{1}$), this condition allows for the operator ${\bm{F}}_{1}$ to be hypomonotone between some pairs of points with a hypomonotonicity parameter that is an absolute constant smaller than one (but can be close to one).
Of course, the condition (31) is required to hold on the boundary of the set ${\mathcal{C}}$ as well for the resolvent to be gradually expansive, and such a condition seems more challenging to ensure. Even so, we note that the current literature addressing root finding problems with operators that are not monotone is only able to address problems under strong conditions either requiring ${\bm{J}}({\bm{x}})$ to be nonexpansive (see [BMW21] for sufficient conditions that ensure the resolvent is nonexpansive) or slightly relaxing this property to the weak MVI condition introduced in [DDJ21]; see, for instance, [AKW24] and references therein.
Finally, another avenue for applying results from this paper to root finding problems is, similar to the example above, by considering problems where ${\bm{F}}$ is the sum of a 1-Lipschitz operator ${\bm{F}}_{1}$ and the normal cone of a closed convex set ${\bm{C}}$ with a nonempty interior and seeking a fixed point of the projection operator ${\bm{T}}({\bm{x}})=\mathrm{Proj}_{{\mathcal{C}}}({\bm{x}}-{\bm{F}}_{1}({\bm{x}}))$. Note that, by standard arguments, points ${\bm{x}}\in{\mathcal{C}}$ with $\left\lVert{\bm{T}}({\bm{x}})-{\bm{x}}\right\rVert\leq\epsilon$ translate into ${\mathcal{O}}(\epsilon)$-approximate roots of ${\bm{F}}.$ In this case, the operator ${\bm{\mathrm{Id}}}-{\bm{T}}$ is also known as the “natural residual” operator (see, e.g., [FP03]). The $\alpha$-gradual nonexpansiveness in this case becomes the following condition for ${\bm{x}},{\bm{y}}\in{\mathcal{C}}$:
$\left\lVert\mathrm{Proj}_{{\mathcal{C}}}({\bm{x}}-{\bm{F}}_{1}({\bm{x}}))-\mathrm{Proj}_{{\mathcal{C}}}({\bm{y}}-{\bm{F}}_{1}({\bm{y}}))\right\rVert$ (33)
$\quad\quad\leq\Big{(}1+\frac{\alpha\max\{\left\lVert\mathrm{Proj}_{{\mathcal{C}}}({\bm{x}}-{\bm{F}}_{1}({\bm{x}}))-{\bm{x}}\right\rVert,\,\left\lVert\mathrm{Proj}_{{\mathcal{C}}}({\bm{y}}-{\bm{F}}_{1}({\bm{y}}))-{\bm{y}}\right\rVert\}}{D}\Big{)}\left\lVert{\bm{x}}-{\bm{y}}\right\rVert.$
This seems like a more tractable condition that may be possible to verify (at least numerically) for different example problems arising in machine learning, such as, for example, those discussed in [DFG20, Das22].
5 Discussion and Numerical Examples
We conclude the paper with a discussion of simple numerical examples and some interesting open problems.
5.1 Numerical Examples
In this subsection, we provide simple numerical examples illustrating the different properties of GHAL and AdaGHAL argued in Section 3. We note here that the purpose of these examples is solely to illustrate the theoretical concepts; more substantive numerical experiments are beyond the scope of the present work.
5.1.1 Adaptation to the Contraction Factor for Nonexpansive Operators
Up to absolute constants, GHAL (Algorithm 1) and AdaGHAL (Algorithm 2) simultaneously recover optimal bounds for both contractive and nonexpansive operators [PR22, DW22]. Observe that the classical Picard iteration—which is optimal for contractive operators—does not converge in general on nonexpansive operators (take, for example, ${\bm{T}}$ to be a rotation operator). On the other hand, classical Halpern iteration with step size $\lambda_{k}=\Theta(\frac{1}{k})$—which is optimal for nonexpansive operators—would not in general converge at a rate faster than $1/k$ even if the operator turned out to be contractive. Consider, for example, the Halpern iteration with step size $\lambda_{k}=\frac{1}{k+1}$ initialized at $x_{0}=1$ and applied to the univariate operator $T(x)=0$. It is not hard to see that each iterate $x_{k}$ of the Halpern iteration would in this case satisfy $x_{k}=\frac{1}{k+1}$ (thus the fixed point error would be $\frac{1}{k+1}$), while the only fixed point is $x_{*}=0.$
Thus, an appealing aspect of Algorithm 2 is that, unlike prior standard iterations, it can converge at an optimal rate without the knowledge of whether the considered operator is contractive or only nonexpansive. This is illustrated using the hard instance from [PR22], which was used in proving oracle lower bounds for finding fixed points of $\gamma$-contractive operators with $\gamma\in(0,1].$ The operator ${\bm{T}}$ is in this case defined by ${\bm{T}}={\bm{R}}_{\gamma},$ where the $i^{\text{th}}$ coordinate of ${\bm{R}}_{\gamma}:{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$, ${\bm{R}}^{(i)}_{\gamma}$, is defined by
${\bm{R}}_{\gamma}^{(i)}({\bm{x}}):=\begin{cases}s-\gamma{\bm{x}}^{(d)},&\text{if }i=1,\\ \gamma{\bm{x}}^{(d-1)},&\text{ if }i\in\{2,\dots,d\}\end{cases},$ (34)
where $s$ is a shift parameter that only affects the initial distance to the fixed point (set to $s=2$ for $\gamma<1$ and to $s=2/\sqrt{d}$ for $\gamma=1$). Figure 1 compares Algorithm 2 (AdaGHAL) to Picard iteration, standard Halpern iteration, and a version of Halpern iteration that employs adaptive restart each time the fixed-point error is halved (see the discussion in Section 3.3). As can be observed from Figure 1, AdaGHAL gracefully adapts to the value of $\gamma$ and closely follows the faster converging among Halpern and Picard iteration, while adaptively restarted Halpern iteration exhibits a noticeably slower linear convergence for contractive operators.
(a) $\gamma=1$
(b) $\gamma=10/11$
(c) $\gamma=5/6$
Figure 1: Comparison of fixed point algorithms for the operator defined by Equation 34 with $d=500$.
(a) $c=0.25$
(b) $c=0.5$
(c) $c=0.75$
Figure 2: Comparison of fixed point algorithms for an operator that is contractive near its fixed point and nonexpansive far from the fixed point.
(a) $c=0.25$
(b) $c=0.5$
(c) $c=0.75$
Figure 3: Comparison of fixed point algorithms for an operator that is nonexpansive near its fixed point and contractive far from the fixed point.
Another appealing property of Algorithm 2 is that it exhibits certain local adaptivity properties. In particular, because the step size is reduced gradually, as the fixed-point error decreases, if on a subset of the algorithm path between subsequent step size updates the operator turns out to be contractive, then the number of iterations between these step size updates would depend on the local contractivity parameter $\gamma.$ We illustrate this property using an operator ${\bm{T}}={\bm{R}}_{1}\circ{\bm{S}}$, where ${\bm{S}}:{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ is defined coordinatewise by ${\bm{S}}^{(i)}({\bm{x}})=f(x_{i})$ and $f:{\mathbb{R}}\to{\mathbb{R}}$ is an even continuous piecewise-linear univariate function with the two defining pieces joining at $|x|=1$. Figure 2 illustrates the results for the case where the pieces $|x|\leq 1$ have the slopes of magnitude 1, while for $|x|>1$ the magnitude is $c\in\{0.25,0.5,0.75\}.$ This corresponds to the case where ${\bm{T}}$ is contractive in a region farther away from the fixed point and only nonexpansive in a region close to the fixed point (for $\|{\bm{x}}\|_{\infty}\leq 1$). As can be observed from Figure 2, AdaGHAL exhibits fast convergence in initial iterations, almost indistinguishable from Picard iteration. Once the algorithm reaches the region where the operator ${\bm{T}}$ is nonexpansive, unlike Picard iteration—which in this case plateaus—AdaGHAL adapts to the slower $1/k$ convergence of Halpern iteration. Although in this second part restarted Halpern iteration gets to a lower error in the nonexpansive region, the convergence rate is similar to the rate of AdaGHAL, while the difference in fixed-point error is by a factor between 4 and 7.
Figure 3 illustrates the results for the case where the pieces $|x|\leq 1$ have the slopes of magnitude $c\in\{0.25,0.5,0.75\}$, while for $|x|>1$ the magnitude is $1$ In other words, in this case the region close to the fixed point is contractive whereas the “further away” region is only nonexpansive. In this case, Picard iteration, restarted Halpern iteration, and AdaGHAL all reach the contractive region within a few iterations and then exhibit fast linear convergence. The standard Halpern iteration on the other hand converges at a slow sublinear $1/k$ rate throughout. AdaGHAL is able to closely follow the convergence of Picard iteration in this case and is much faster than restarted Halpern iteration, similar to the examples shown in the center and right plot in Figure 1.
5.1.2 Convergence for More General Lipschitz Operators
We now consider Lipschitz operators that are not necessarily nonexpansive, to illustrate our results concerning mildly expansive operators. We note here that existing algorithms such as Picard iteration, Halpern iteration, and restarted Halpern iteration, have no theoretical convergence guarantees for such operators.
We consider here a simple example of an operator obtained as a composition of three operators: ${\bm{T}}={\bm{P}}_{\mathcal{B}}\circ{\bm{R}}\circ{\bm{S}}$ defined as follows. ${\bm{P}}_{\mathcal{B}}$ is a projection operator that projects its argument onto the unit $\ell_{2}$ ball: ${\bm{P}}_{\mathcal{B}}({\bm{x}})=\frac{{\bm{x}}}{\max\{1,\|{\bm{x}}\|_{2}\}}.$ Clearly, it is nonexpansive w.r.t. the $\ell_{2}$ norm. ${\bm{R}}$ is equal to the operator ${\bm{R}}_{\gamma^{\prime}}|_{\gamma^{\prime}=1}$ defined in (34), which was already argued to be nonexpansive. Finally, for ${\bm{S}}$, we consider a piecewise-linear operator, defined as follows:
${\bm{S}}^{(i)}({\bm{x}})=\begin{cases}\gamma{\bm{x}}^{(i)},&\text{ if }|{\bm{x}}^{(i)}|\leq 1-c,\\ {\bm{x}}^{(i)}+(\gamma-1)(1-c),&\text{ if }|{\bm{x}}^{(i)}|>1-c,\end{cases}$ (35)
where we choose the values of $c$ to be between zero and one. When $c=1,$ ${\bm{S}}$ is a linear operator with Lipschitz constant equal to $1,$ so it is nonexpansive. When $c=0,$ ${\bm{S}}$ is a linear operator with Lipschitz constant equal to $\gamma.$ For other values of $c$ between zero and one, each coordinate of ${\bm{S}}$ has two linear pieces: one with Lipschitz constant $\gamma$ and one with Lipschitz constant equal to one. It is not hard to see that ${\bm{S}}$ is $\gamma$-Lipschitz (w.r.t. any $\ell_{p}$ norm, and, in particular, w.r.t. the $\ell_{2}$ norm). Since ${\bm{P}}_{\mathcal{B}}$ and ${\bm{R}}$ are nonexpansive and ${\bm{S}}$ is $\gamma$-Lipschitz, ${\bm{T}}={\bm{P}}_{\mathcal{B}}\circ{\bm{R}}\circ{\bm{S}}$ is $\gamma$-Lipschitz.
(a) $c=0.1,$ $\gamma=1+10^{-4}$
(b) $c=0.1,$ $\gamma=1+10^{-3}$
(c) $c=0.1$, $\gamma=1+10^{-2}$
(d) $c=0.5,$ $\gamma=1+10^{-4}$
(e) $c=0.5,$ $\gamma=1+10^{-3}$
(f) $c=0.5$, $\gamma=1+10^{-2}$
(g) $c=0.9,$ $\gamma=1+10^{-4}$
(h) $c=0.9,$ $\gamma=1+10^{-3}$
(i) $c=0.9$, $\gamma=1+10^{-2}$
Figure 4: Comparison of fixed point algorithms for an operator that is $\gamma$-Lipschitz, defined by ${\bm{T}}={\bm{P}}_{\mathcal{B}}\circ{\bm{R}}\circ{\bm{S}},$ with operators ${\bm{P}}_{\mathcal{B}}$ being the projection onto the centered unit $\ell_{2}$ ball, ${\bm{R}}$ defined by (34) with $\gamma=1$, and ${\bm{S}}$ defined by (35). GHAL exhibits an overall more robust convergence to approximate fixed points of ${\bm{T}}$ compared to other algorithms.
We illustrate convergence of Picard iteration, Halpern iteration, restarted Halpern iteration, and GHAL for $c\in\{0.1,0.5,0.9\}$, $\gamma\in 1+\{10^{-2},10^{-3},10^{-4}\},$ and $d=100$ in Figure 4. The parameters of GHAL in all the examples shown in Figure 4 are set to $\beta=0.975,$ $\beta^{\prime}=1/1.01,$ following Theorem 2. We observe from Figure 4 that GHAL exhibits an overall more robust convergence to approximate fixed points of ${\bm{T}}$ than other algorithms. In particular, neither Picard nor standard Halpern iteration are competitive on any of the considered examples, apart from, perhaps, the one shown in the bottom right plot. Instead, both algorithms either exhibit oscillatory behavior (e.g., the first two plots in the right column) or otherwise plateau at a higher fixed-point error than restarted Halpern and GHAL. On the other hand, restarted Halpern iteration in some instances (middle column) converges faster than GHAL. We conjecture that what happens in this case is that restarted Halpern iteration is able to reduce the fixed-point error quickly for large values of the step size $\lambda_{k}$, correcting for $\gamma>1,$ before reaching a restart condition and again working with large step sizes. We see the limitations of this strategy as the value of $\gamma$ is increased (right column in Figure 4, top two rows), as restarted Halpern iteration, similar to standard Halpern iteration, exhibits oscillatory behavior. GHAL however stabilizes at a lower value of the fixed-point error in this case.
5.2 Further Discussion and Open Problems
While the present work provides nontrivial classes of problems for which fixed points can be efficiently approximated, there are several avenues for future research that would deepen our understanding of this area and possibly resolve some of the open questions in game theory such as those discussed in [Das22].
5.2.1 Improving Lower Bounds (or Reducing the Gap between Upper and Lower Bounds)
As discussed in Section 2.2, the classical lower bound for solving fixed-point equations with $\gamma$-Lipschitz operators mapping $[0,D]^{d}$ ($d\geq 3$) to itself, due to [HPV89], bounds below the worst case number of required queries to the operator to find a point ${\bm{x}}\in[0,D]^{d}$ such that $\left\lVert{\bm{T}}({\bm{x}})-{\bm{x}}\right\rVert\leq\epsilon$, where $\epsilon\in(0,1/10),$ by
$\Big{(}c\Big{(}\frac{1}{\epsilon}-10\Big{)}(\gamma-1)D\Big{)}^{d-2},$
where $c\geq 10^{-5}.$ The authors of [HPV89] conjectured that $c$ can in fact be made close to one. Obtaining such a lower bound with a larger value of $c$ (ideally, $c=1$) and removing the shift by 10 in the statement of the lower bound would effectively establish that our results get to the boundary of what is information-theoretically possible for polynomial-time algorithms. Note here that in such a statement of a lower bound, one would need to be careful with the interval for $\epsilon,$ as any point in the feasible set has fixed-point error bounded by $D.$ We also note here that none of the recent lower bounds such as [ADD+25] implies a tighter range of constants for related settings, as obtaining tighter constants in the lower bounds was not their focus.
Further, if, say, a lower bound of the form $\big{(}\frac{\gamma-1}{\epsilon}D\big{)}^{d}$ were to be obtained for $\gamma$ in any subset of $(1,2)$ and $\epsilon\in(0,\gamma-1)$, by embedding the $d$-dimensional problem in an infinite-dimensional Banach space and increasing the dimension, this would imply the impossibility of solving “minimal displacement problem” to error $1-1/\gamma+{\overline{\epsilon}}$ for small ${\overline{\epsilon}}>0,$ as was discussed in Section 2.2. As a result, it would also disprove the conjecture of [BB96] about convergence of the KM iteration in Banach spaces for $\gamma$-Lipschitz operators.
Conversely, any improvement to the obtained results on approximating $\left\lVert{\bm{T}}({\bm{x}})-{\bm{x}}\right\rVert$ that nontrivially reduces the gap between upper and lower bounds would be interesting.
5.2.2 Convergence Under Local Error Bounds
As discussed earlier, conditions on expansive operators ${\bm{T}}$ that enforce an upper bound on some notion of a distance between an element ${\bm{x}}$ and the set of fixed points of ${\bm{T}}$ as a function of the fixed point error $\left\lVert{\bm{T}}({\bm{x}})-{\bm{x}}\right\rVert$ has been used in prior work [Nat93, LTT18, LL21, HLS23] to ensure convergence to solutions with arbitrarily small fixed-point error is possible, often at a (locally) linear rate (though the “constant” of linear convergence can be quite—even exponentially in the dimension—small). Such results rely upon the expansion of ${\bm{T}}$ being controlled by the parameter of the enforced local error bound. It appears possible (with some technical work) to generalize our results to settings where such a local error bound applies, possibly by slightly modifying the introduced algorithms. Some questions that we find interesting in this context are: Would the range of problems solvable by such algorithms to an arbitrarily small error be extended in some nontrivial way? Even without requiring convergence to an arbitrarily small error, can we improve the fixed-point error bound of the algorithm? How are these two notions/requirements intertwined?
5.2.3 Gradual Expansion in Applications
Finally, this work introduced the class of gradually expansive operators without providing concrete examples for which standard fixed-point iterations would fail. Section 4.1 provides an example of an operator that is $\alpha$-gradually expansive (but not nonexpansive) and proves additional sufficient conditions for constructing such operators. It would be interesting to construct other nontrivial examples of operators that are $\alpha$-gradually expansive and/or prove additional structural properties for the associated class of fixed-point problems.
In a different direction, proving gradual expansiveness for operators arising in common game theoretic settings whose complexity is not yet understood (or even verifying it numerically) could help address some of the recognized open problems in game theory and machine learning, such as those outlined in [Das22].
Acknowledgments
This work was supported in part by an AFOSR Young Investigator Program award, under the contract number FA9550-24-1-0076 and by the U.S. Office of Naval Research under contract number N00014-22-1-2348. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U.S. Department of Defense.
The author thanks Heinz Bauschke, Roberto Cominetti, and Russell Luke for useful comments and pointers to the relevant literature.